### v3

# On the Implicit Relation Between Low-Rank Adaptation and Differential Privacy

Abstract
A significant approach in natural language processing involves large-scale pre-training models on general domain data followed by their adaptation to specific tasks or domains. As models grow in size, full fine-tuning all of their parameters becomes increasingly impractical. To address this, some methods for low-rank task adaptation of language models have been proposed, e.g., LoRA and FLoRA. These methods keep the pre-trained model weights fixed and incorporate trainable low-rank decomposition matrices into some layers of the transformer architecture, called adapters. This approach significantly reduces the number of trainable parameters required for downstream tasks compared to full fine-tuning all parameters. In this work, we look at low-rank adaptation from the lens of data privacy. We show theoretically that the low-rank adaptation used in LoRA and FLoRA is equivalent to injecting some random noise into the batch gradients w.r.t the adapter parameters, and we quantify the variance of the injected noise. By establishing a Berry-Esseen type bound on the total variation distance between distribution of the injected noise and a Gaussian distribution with the same variance, we show that the dynamics of low-rank adaptation is close to that of differentially private fine-tuning of the adapters. Finally, using Johnson-Lindenstrauss lemma, we show that when augmented with gradient scaling, low-rank adaptation is very close to performing DPSGD algorithm with a fixed noise scale to fine-tune the adapters. These theoretical findings suggest that unlike other existing fine-tuning algorithms, low-rank adaptation provides privacy w.r.t the fine-tuning data implicitly.
1 Introduction
Stochastic Gradient Descent (SGD) is the power engine of training deep neural networks, which updates parameters of a model by using a noisy estimation of the gradient. Modern deep learning models, e.g., GPT-3 (Brown et al., 2020) and Stable Diffusion (Rombach et al., 2022), have a large number of parameters, which induces a large space complexity for their training with SGD. Using more advanced methods, which track various gradient statistics to stabilize and accelerate training, exacerbates this space complexity (Duchi et al., 2011). For instance, momentum technique reduces variance by using an exponential moving average of gradients (Cutkosky and Orabona, 2019). Also, gradient accumulation (Wang et al., 2013) reduces variance by computing the average of gradients in the last few batches, which simulates a larger effective batch size. Both the methods suffer from an extra space complexity during training/fine-tuning time.
Some works try to reduce the space complexity of fine-tuning large models by tuning a subset of parameters and storing the information about only a portion of them (Houlsby et al., 2019; Ben Zaken et al., 2022). LoRA (Hu et al., 2021) is such an algorithm, which only updates some of the parameter matrices, called adapters, by restricting their updates to be a low-rank matrix. This low-rank restriction considerably reduces the number of trainable parameters, at the cost of limiting the optimization space of the adapter parameters. Another parameter-efficient training technique, called ReLoRA (Lialin et al., 2023), utilizes low-rank updates to train high-rank networks to eliminate the constraint of LoRA mentioned above. Similarly, the work in (Hao et al., 2024) identifies that the dynamics of LoRA can be approximated by a random matrix projection. Based on this interesting finding, the work proposes FLoRA to achieve high-rank updates by resampling the random projection matrices, which enables achieving higher utility while still enjoying the sublinear space complexity of LoRA.
On the other hand, from the lens of data privacy, the fine-tuning data often happens to be privacy sensitive. In such scenarios, Differentially Private (DP) fine-tuning algorithms have been used to provide rigorous privacy guarantees w.r.t the data. DP full fine-tuning runs DPSGD (Abadi et al., 2016) on the the fine-tuning data to update all the existing parameters in a model. However, due to computing gradients and clipping them for every data sample, running DPSGD on all parameters induces high space complexities, even higher than non-private full fine-tuning of all parameters.
In this work, we draw a connection between low-rank adaptation and differentially private fine-tuning of the adapters. We show that the random projection existing in the dynamics of LoRA/FLoRA is equivalent to injecting some random noise into the batch gradients w.r.t the adapters, which is very close to what DPSGD does for fine-tuning adapters privately. We also quantify the variance of the injected noise as a function of the rank of adaptation, and show that it increases as the rank of adaptation decreases: the smaller the rank of adaptation, the larger the variance of the injected noise. Furthermore, in order to evaluate the closeness of this injected noise to Gaussian noise with the same variance, we bound the total variation (TV) distance between the distribution of the injected noise and the pure Gaussian noise used in DPSGD and show that this bound (dissimilarity) decreases as the rank used in LoRA/FLoRA increases. Our derivations suggest that, although not being exactly the same, low-rank adaptation and DP fine-tuning of the adapters are very close to each other in terms of their dynamics. This implies that, besides reducing the space complexity for task adaptation of language models, low rank adaptation can provide privacy w.r.t the fine-tuning data implicitly without inducing the high space complexity of running DPSGD on adapters or all parameters.
The highlights of our contributions are the followings:
  • We show that low-rank adaptation with LoRA/FLoRA is equivalent to injection of some random noise into the batch gradients w.r.t the adapters (section 3).
  • We find the variance of the noise injected into the adapters’ gradients, and show that the injected noise approaches a Gaussian distribution as the input dimension of the adaptation layer increases (lemma 3.1).
  • We bound the total variation distance between the distribution of the injected noise and a pure Gaussian noise with the same variance. The bound decreases as the input dimension of the adaptation layer and the adaptation rank increase (lemma 4.2).
  • Finally, we show that the dynamics of low-rank adaptation is close to DP fine-tuning of the adapters, and this similarity increases when low-rank adaptation is augmented with gradient scaling. This implies an implicit connection between low-rank adaptation and DPSGD. Indeed, the former is very close to DPSGD with a fixed noise scale, which depends on the adaptation rank, the batch size used during adaptation and the output dimension of the adaptation layer (section 5).
2 Dynamics of Low-Rank Task Adaptation
We start by studying the dynamics of low-rank adaptation (Hao et al., 2024). In order to update a pre-trained adapter weight $W\in\mathbb{R}^{n\times m}$, LoRA incorporates low-rank decomposition matrices $B\in\mathbb{R}^{n\times r}$ and $A\in\mathbb{R}^{r\times m}$, where $r\ll\min\{n,m\}$, and performs the forward pass in a layer as:
$y=(W+BA)x=Wx+BAx,$ (1)
Figure 1: Low-rank decomposition of LoRA/FLoRA for task adaptation.
where $x\in\mathbb{R}^{m}$ is the input of the current layer and $y\in\mathbb{R}^{n}$ is the pre-activation output of the current layer (see fig. 1). It is common to initialize $B$ with an all-zero matrix and $A$ with a normal distribution. More specifically, the entries of $A$ are sampled from $\mathcal{N}(0,\sigma^{2})$ with $\sigma^{2}=\frac{1}{r}$. When back-propagating, gradient of the used loss function $\mathcal{L}$ w.r.t the matrix $W$ is
$\nabla_{W}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial y}\cdot\frac{\partial y}{\partial W}=\frac{\partial\mathcal{L}}{\partial y}\cdot x^{\top},$ (2)
where $\frac{\partial\mathcal{L}}{\partial y}\in\mathbb{R}^{n\times 1}$ and $x^{\top}\in\mathbb{R}^{1\times m}$. LoRA calculates the gradients w.r.t only A and B, which can be found as follows:
$\frac{\partial\mathcal{L}}{\partial A}=\frac{\partial BA}{\partial A}\cdot\frac{\partial\mathcal{L}}{\partial BA}=B^{\top}\cdot\frac{\partial\mathcal{L}}{\partial y}\cdot\frac{\partial y}{\partial BA}=B^{\top}\cdot\frac{\partial\mathcal{L}}{\partial y}\cdot x^{\top}=B^{\top}(\nabla_{W}\mathcal{L}).$ (3)
Similarly,
$\frac{\partial\mathcal{L}}{\partial B}=\frac{\partial\mathcal{L}}{\partial BA}\cdot\frac{\partial BA}{\partial B}=\frac{\partial\mathcal{L}}{\partial y}\cdot\frac{\partial y}{\partial BA}\cdot A^{\top}=\frac{\partial\mathcal{L}}{\partial y}\cdot x^{\top}\cdot A^{\top}=(\nabla_{W}\mathcal{L})A^{\top}.$ (4)
Hence, $\frac{\partial\mathcal{L}}{\partial A}\in\mathbb{R}^{r\times m}$ and $\frac{\partial\mathcal{L}}{\partial B}\in\mathbb{R}^{n\times r}$. As observed in eq. 3 and eq. 4, LoRA down-projects the batch gradient $\nabla_{W}\mathcal{L}$ from $\mathbb{R}^{n\times m}$ to a lower dimension, and updates the matrices $A$ and $B$ with the resulting projections of $\nabla_{W}\mathcal{L}$. In fact, it was found in (Hao et al., 2024) that LoRA recovers the well-known random projection method (Dasgupta, 2000; Bingham and Mannila, 2001). We restate the following thorem from (Hao et al., 2024) without restating the proof:
Theorem 2.1 ((Hao et al., 2024), Theorem 2.1).
Let LoRA update matrices $A$ and $B$ with SGD:
$A^{t+1}\leftarrow A^{t}-\eta\frac{\partial\mathcal{L}}{\partial A^{t}}=A^{t}-\eta B^{t}{{}^{\top}}(\nabla_{W}\mathcal{L}^{t}),$ (5)
$B^{t+1}\leftarrow B^{t}-\eta\frac{\partial\mathcal{L}}{\partial B^{t}}=B^{t}-\eta(\nabla_{W}\mathcal{L}^{t})A^{t}{{}^{\top}},$ (6)
where $\eta$ is the learning rate. We assume $\|\sum_{t=0}^{T}\nabla_{W}\mathcal{L}^{t}\|_{F}\leq L$ for every $T$, which implies that the model stays within a finite Euclidean ball. In this case, the dynamics of $A^{t}$ and $B^{t}$ are given by
$A^{t}=A^{0}+\eta A^{0}f_{A}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ B^{t}=\eta f_{B}(t)A^{0}{{}^{\top}},$ (7)
where the forms of $f_{A}(t)\in\mathbb{R}^{m\times m}$ and $f_{B}(t)\in\mathbb{R}^{n\times m}$ are expressed in the proof. In particular, $\|f_{A}(t)\|_{2}\leq\frac{\eta L^{2}\big{(}1-(\eta^{2}L^{2})^{t}\big{)}}{1-\eta^{2}L^{2}}$ for every $t$.
Let’s denote the total changes of $A$ and $B$ after $T$ steps as $\Delta A$ and $\Delta B$, respectively. Then, the forward pass in eq. 1 changes to:
$\big{(}W+(B^{0}+\Delta B)(A^{0}+\Delta A)\big{)}x=\big{(}W+\Delta BA^{0}+\Delta B\Delta A\big{)}x,$ (8)
where we have substituted $B^{0}=\mathbf{0}\in\mathbb{R}^{n\times r}$. From eq. 7 and substituting the values of $\Delta A$ and $\Delta B$ after $T$ rounds of updating $A$ and $B$, we have:
$W+\Delta BA^{0}+\Delta B\Delta A=W+\eta f_{B}(T)A^{0}{{}^{\top}}A^{0}+\eta^{2}f_{B}(T)A^{0}{{}^{\top}}A^{0}f_{A}(T).$ (9)
Also, from theorem 2.1, we have $\|f_{A}(T)\|_{2}\leq\|f_{A}(T)\|_{F}\leq\frac{\eta L^{2}\big{(}1-(\eta^{2}L^{2})^{T}\big{)}}{1-\eta^{2}L^{2}}$. Hence, if $\eta\ll 1/L$, we have $\lim_{T\to\infty}\eta\|f_{A}(T)\|_{2}=\lim_{T\to\infty}\frac{(\eta L)^{2}\big{(}1-(\eta L)^{(2T)}\big{)}}{1-(\eta L)^{2}}\ll 1$. Therefore, the last term in eq. 9 is significantly smaller than the second term. Hence, the second term dominates the final update weight. Therefore, as suggested in (Hao et al., 2024) and confirmed with their experimental results, we can closely approximate LoRA by freezing $A$ at its initialized value $A^{0}$ and training only the matrix $B$. In this case,
$W+\Delta BA^{0}+\Delta B\Delta A$ $=W+\Delta BA^{0}=W+\eta\tilde{f}_{B}(T)A^{0}{{}^{\top}}A^{0},$ (10)
where $\tilde{f}_{B}(0)=\mathbf{0}$ and $\tilde{f}_{B}(t+1)=\tilde{f}_{B}(t)-\nabla_{W}\mathcal{L}^{t}$. Equivalently, $\tilde{f}_{B}(T)=-\sum_{t=0}^{T-1}\nabla_{W}\mathcal{L}^{t}$. Substituting this into the equation above, we get:
$W+\Delta BA^{0}+\Delta B\Delta A=W+\Delta BA^{0}=W-\eta\sum_{t=0}^{T-1}\big{[}(\nabla_{W}\mathcal{L}^{t})A^{0}{{}^{\top}}A^{0}\big{]},$ (11)
where the last term shows the exact parameter change after $T$ rounds of performing SGD on the adapter matrix $B$. Therefore, low rank adaptation with LoRA can be viewed as performing a random projection of stochastic batch gradient $\nabla_{W}\mathcal{L}^{t}$ in every step $t$ by matrix $A^{0}{{}^{\top}}$ and projecting it back by matrix $A^{0}$. FLoRA (Hao et al., 2024) proposes to resample the random matrix $A^{0}$ at each step to get a high rank update $\Delta B$ for the matrix $B$. Hence, FLoRA can also be viewed as performing a random projection of stochastic batch gradient $\nabla_{W}\mathcal{L}^{t}$ in every step $t$ by a different random matrix $A{{}^{\top}}$ and projecting it back by its transpose.
Having understood the connection between low-rank adaptation in LoRA/FLoRA and random projection, in the next section, we show that this random projection and back projection performed in each time step is equivalent to adding some random noise to each element of $\nabla_{W}\mathcal{L}^{t}$. This is our first step towards establishing the connection between low-rank adaptation and differential privacy.
3 Random Noise Injected by Low-Rank Adaptation
In this section, we present our analysis based on LoRA, which employs a fixed projection matrix $A^{0}$. Our analysis holds for various LoRA variants, including FLoRA. As illustrated in section 3, the parameter update after $T$ rounds of stochastic gradient descent (SGD) is given by:
$W+\Delta BA^{0}+\Delta B\Delta A$ $=W-\eta\sum_{t=0}^{T-1}\bigg{[}(\nabla_{W}\mathcal{L}^{t})A^{0}{{}^{\top}}A^{0}\bigg{]}$
$=W-\eta\sum_{t=0}^{T-1}\bigg{[}\underbrace{\nabla_{W}\mathcal{L}^{t}}_{\textit{ fine-tuning adapters}}\underbrace{-\nabla_{W}\mathcal{L}^{t}(A^{0}{{}^{\top}}A^{0}-\mathbb{I}_{m})}_{\textit{noise}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathbb{R}^{n\times m}}\bigg{]},$ (12)
The first term in the sum represents the batch gradient that would be obtained through fine-tuning the adapter $W$ using SGD. The second term represents the noise introduced by the low-rank adaptation. Thus, the low-rank adaptation introduces noise to each batch gradient $\nabla_{W}\mathcal{L}^{t}$, and the gradient step is taken with this noisy gradient. We are now particularly interested in the behavior of this noise term, which is added to each batch gradient $\nabla_{W}\mathcal{L}^{t}$ in every step $t$. Recall that the entries of $A^{0}$ were sampled from $\mathcal{N}(0,\frac{1}{r})$ (see fig. 1), and that each of the $r$ columns of $A^{0}{{}^{\top}}$ is an $m$-dimensional Gaussian random variable. Consequently, $A^{0}{{}^{\top}}A^{0}$ follows a Wishart distribution with $r$ degrees of freedom (Bhattacharya and Burman, 2016), which is the multivariate generalization of the chi-squared distribution. Therefore, for any $q\in\mathbb{R}^{1\times m}$, $q\cdot(A_{0}^{\top}A_{0}-\mathbb{I}_{m})$ is a weighted sum of multiple chi-squared random variables, which implies that the result follows a Gaussian distribution approximately, according to the Central Limit Theorem (CLT) (Bhattacharya et al., 2016). We prove the following lemma concerning the noise term in section 3.
Lemma 3.1.
Let $A\in\mathbb{R}^{r\times m}$ be a matrix with i.i.d entries sampled from $\mathcal{N}(0,\frac{1}{r})$. Given a fixed $q\in\mathbb{R}^{1\times m}$, the distributions of elements of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ approach the Gaussian distribution $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$, as $m$ increases.
The result above can be extended to matrices multiplication, as in section 3: for a matrix $Q\in\mathbb{R}^{n\times m}$ and as $m$ grows, the product $G=Q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{n\times m}$ approaches a Gaussian distribution, where $G_{i,j}$ ($1\leq i\leq n$) has distribution $\mathcal{N}(0,\frac{\|[Q]_{i,:}\|^{2}}{r})$, where $[Q]_{i,:}$ is the $i$-th row of $Q$. The lemma above shows that the last term in section 3 can indeed be looked at as a random noise term. As a weak assumption, hereafter, we assume that all the rows of $\nabla_{w}\mathcal{L}^{t}$ have a non-zero norm, which results in all the elements in the noise term having a variance greater than zero.
Although lemma 3.1 was proved for when $m$ approaches infinity, in practical scenarios $m$ is limited. Hence, the distribution of the injected noise in not pure Gaussian. In the next section, we bound the deviation of the noise distribution from a pure Gaussian distribution.
4 Bounding the Distance to the Normal Law
Despite having proved lemma 3.1 when $m$ approaches infinity, yet we need to quantify the distance between the distribution of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ to the bona fide Gaussian distribution for limited values of $m$ in practical scenarios. In this section, we derive a closed form upper-bound for the total variation distance between the distribution of each element of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ and the Gaussian distribution with the same mean and variance.
Consistent with the notations in Theorem A.4 in the appendix, suppose $X_{1},\ldots,X_{n}$ are $n$ independent random variables with $\mathbb{E}[X_{i}]=0$ (mean) and $\texttt{Var}[X_{i}]=\sigma_{i}^{2}>0$ (variance). Define $S_{n}=\sum_{i=1}^{n}X_{i}$ and let $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$. Assuming $Z_{n}=\frac{S_{n}}{s_{n}}$, and having Lindeberg’s condition satisfied (see theorem A.3 and theorem A.4 in the appendix), the normalized sum $Z_{n}$ has standard normal distribution in a weak sense for a bounded $n$. More precisely, the closeness of the cumulative distribution function (CDF) $F_{n}(x)=\texttt{Pr}\{Z_{n}\leq x\}$ to the standard normal CDF
[…]
has been studied intensively in terms of the Lyapounov ratios
$L_{t}=\frac{\sum_{i=1}^{n}\mathbb{E}[|X_{i}|^{t}]}{s_{n}^{t}}.$ (14)
Particularly, if all $X_{i}$ have a finite third absolute moment $\mathbb{E}[|X_{i}|^{3}]$, the classical Berry-Esseen theorem (gustav Esseen, 1945; Feller, 1971; Petrov, 1975) bounds the Kolmogrov distance between $F_{n}(x)$ and $\Phi(x)$:
$\sup_{x}|F_{n}(x)-\Phi(x)|\leq CL_{3},$ (15)
where $C$ is an absolute constant. In the more general case of sum of independent random variables (and not necessarily i.i.d random variables), which we are interested in, the number of summand variables $n$ implicitly affects the value of $L_{3}$, and for this case, the work in (Bobkov et al., 2011) bounds the difference between $F_{n}(x)$ and $\Phi(x)$ in terms of generally stronger distances of total variation and entropic distance. Considering the $X_{i}$ above, let $D(X_{i})$ denote the KL divergence between distribution of $X_{i}$ and Gaussian distribution $\mathcal{N}(0,\sigma_{i}^{2})$, i.e., the KL divergence between $X_{i}$ and a Gaussian with the same variance. We have the following theorem about the total variation distance between $F_{n}$ and $\Phi$:
Theorem 4.1 ((Bobkov et al., 2011), theorem 1.1).
Assume that the independent random variables $X_{1},\ldots,X_{n}$ have finite third absolute moments, and that $D(X_{i})\leq D$, where $D$ is a non-negative number. Then,
$\|F_{n}(x)-\Phi(x)\|_{TV}\leq C_{D}L_{3},$ (16)
[…]
Having the theorem above, we can now derive a Berry-Esseen type bound for the total variation distance between each element of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ in lemma 3.1 and the normal law $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$: we need to find the third Lyapounov ratio for the summands contributing to each element, as in eq. 16. To this end, we state and prove the following lemma:
Lemma 4.2.
Let $A\in\mathbb{R}^{r\times m}$ be a matrix with i.i.d entries sampled from $\mathcal{N}(0,\frac{1}{r})$. Given a fixed $q\in\mathbb{R}^{1\times m}$ with elements $0<c\leq|q_{i}|\leq C$, let $u=q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$. Let $u_{i}$ be the $i$-th element of $u$ and $Q_{m}(x)=\texttt{Pr}\{u_{i}\leq x\}$. Also, let $\Phi(x)$ be the CDF of $z\sim\mathcal{N}(0,\frac{\|q\|^{2}}{r})$. Then:
$\|Q_{m}(x)-\Phi(x)\|_{TV}\in\mathcal{O}\bigg{(}\frac{1}{\sqrt{mr}}\bigg{)}.$ (17)
The lemma above states that the distribution of each element of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ approaches to $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$ with rate $\frac{1}{\sqrt{mr}}$ in terms of their total variation distance. This result shows the elements of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ approach to Gaussian $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$ as $m$ and $r$ increase. Having the interesting result above, we can now benefit from the useful coupling characterization of the total variation distance to establish a more understandable relation between each element of the product above and the Gaussian distribution $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$.
The coupling characterization of the total variation distance.
For two distributions $P$ and $Q$, a pair of random variables $(X,Y)$, which are defined on the same probability space, is called a coupling for $P$ and $Q$ if $X\sim P$ and $Y\sim Q$ (Levin et al., 2008; Devroye et al., 2023). A very useful property of total variation distance is the coupling characterization (see proposition 4.7 in (Levin et al., 2008)):
$\|P-Q\|_{TV}\leq t$ if and only if there exists a coupling $(X,Y)$ for them such that $\texttt{Pr}\{X\neq Y\}\leq t$.
Hence, we can use the coupling characterization above and get to the following lemma directly from from lemma 4.2.
Lemma 4.3.
Let $A\in\mathbb{R}^{r\times m}$ be a matrix with i.i.d entries sampled from $\mathcal{N}(0,\frac{1}{r})$. Given a fixed $q\in\mathbb{R}^{1\times m}$ with elements $0<c\leq|q_{i}|\leq C$, let $u=q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$. Let $u_{i}$ be the $i$-th element of $u$. Then there exists a coupling $(u_{i},z)$, where $z\sim\mathcal{N}(0,\frac{\|q\|^{2}}{r})$ and
$\texttt{Pr}\{u_{i}\neq z\}\in\mathcal{O}\bigg{(}\frac{1}{\sqrt{mr}}\bigg{)}.$ (18)
The lemma above means that each element $u_{i}$ follows a mixture of distributions: $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$ with weight $w_{g}$ and another distribution $M$, which we dont know, with weight $(1-w_{g})\in\mathcal{O}\big{(}\frac{1}{\sqrt{mr}}\big{)}$. The larger $mr$, the closer the mixture distribution gets to pure Gaussian distribution $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$. Having the results above, we can now draw a clear connection between low-rank adaptation and DP.
5 Connecting Low Rank Adaptation to DP with Gradient Scaling
Based on section 3 and our understandings from lemma 4.3, low rank adaptation (with rank $r$) of adapter parameter $W\in\mathbb{R}^{n\times m}$ at time step $t$ is equivalent to fine-tuning it with the noisy stochastic batch gradients ~$\tilde{\nabla}_{W}\mathcal{L}^{t}=\nabla_{W}\mathcal{L}^{t}+N^{t}$, where $N^{t}\in\mathbb{R}^{n\times m}$ is a noise-term with Gaussian-like distribution: $\texttt{Pr}\{N_{i,j}^{t}\neq z_{i}^{t}\}\in\mathcal{O}\big{(}\frac{1}{\sqrt{mr}}\big{)}$, where $z_{i}^{t}\sim\mathcal{N}(0,\frac{\|[\nabla_{W}\mathcal{L}^{t}]_{i,:}\|^{2}}{r})$. In other words, low-rank adaptation adds noise to each row of batch gradient $\nabla_{W}\mathcal{L}^{t}$, and the standard deviation of the noise added to the elements of the row $i$ is proportional to the $\ell_{2}$ norm of row $i$. Also, as $mr$ grows, i.e., the input dimension of the adaptation layer ($m$) increases or the adaptation rank increases ($r<m$), the distribution of noise element $N_{i,j}^{t}$ gets closer to $\mathcal{N}(0,\frac{\|[\nabla_{W}\mathcal{L}^{t}]_{i,:}\|^{2}}{r})$. This operation is very similar to what DPSGD (Abadi et al., 2016) does for adding noise to each element of the clipped batch gradients w.r.t the adapter parameters: at the $t$-th gradient update step on a current adapter parameter $W$, DPSGD computes the following noisy batch gradient on a batch $\mathcal{B}^{t}$ with size $b$:
~$\tilde{\nabla}_{W}\mathcal{L}^{t}=\frac{1}{b}\bigg{[}\Big{(}\sum_{i\in\mathcal{B}^{t}}\bar{\nabla}_{W}\mathcal{L}_{i}^{t}\Big{)}+\mathcal{N}(0,\sigma_{\texttt{DP}}^{2})\bigg{]},$ (19)
where ¯$\bar{\nabla}_{W}\mathcal{L}_{i}^{t}=\texttt{clip}(\nabla_{W}\mathcal{L}_{i}^{t},c)$, $c$ is a clipping threshold, and $\mathcal{B}^{t}$ is the batch of samples at time step $t$. Also, $\sigma_{\texttt{DP}}=c\cdot z$, where $z$ is the noise scale determining the resulting privacy guaranty parameters. The main difference between the noise addition mechanism in low-rank adaptation (section 3) and that in DPSGD (eq. 19) is that DPSGD adds a noise with a fixed variance $\sigma_{\texttt{DP}}^{2}$ to all elements of the clipped batch gradient, and also there is no sample gradient clipping happening in low rank adaptation. In the following, we show that how this clipping can be introduced in low rank adaptation with almost no cost by using Johnson-Lindenstrauss Lemma. This also leads to the same noise variance for all elements. We first state a version of the lemma in the following.
Theorem 5.1 ((Matousek, 2008), Theorem 3.1).
Let $m$ be an integer, $\Delta\in(0,\frac{1}{2}]$, and $p\in(0,1)$. Also, let $r=\Delta^{-2}log(\frac{2}{p})$. Let us define a random linear map $T:\mathbb{R}^{m}\to\mathbb{R}^{r}$ by
$T(x)_{i}=\frac{1}{\sqrt{r}}\sum_{j=1}^{m}R_{ij}x_{j},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ i=1,\cdots,r$ (20)
where the $R_{ij}$ are independent standard normal variables. Then, for every $x\in\mathbb{R}^{m}$, we have:
$\texttt{Pr}[\|T(x)\|\geq(1+\Delta)\|x\|]\leq\frac{p}{2},$
$\texttt{Pr}[\|T(x)\|\leq(1-\Delta)\|x\|]\leq\frac{p}{2}.$ (21)
Also, as a direct consequence of the above inequalities:
$\texttt{Pr}\bigg{[}\frac{\|T(x)\|}{(1+\Delta)}\leq\|x\|\leq\frac{\|T(x)\|}{(1-\Delta)}\bigg{]}\geq 1-p.$ (22)
The theorem above directly relates to the random projection mapping $A^{\top}$ observed in LoRA/FLoRA: let us define the mapping $T$ in theorem 5.1 to be $T(x)=xA^{\top}$. Then, we know that for a sample $i$ in a batch of samples with size $b$, $\nabla_{B^{t}}\mathcal{L}_{i}^{t}=T(\nabla_{W^{t}}\mathcal{L}_{i}^{t})$. More precisely, let’s fix $\Delta\in(0,\frac{1}{2}]$ and define $p$ implicitly with $r=\Delta^{-2}\log(\frac{2}{p})$ ($r$ is the adaptation rank). Then, according to eq. 22, for every sample $i$ in a batch $\mathcal{B}^{t}$ and every row $l\in[1,n]$, we have:
$\big{\|}[\nabla_{B^{t}}\mathcal{L}_{i}^{t}]_{l,:}\big{\|}=(1-\Delta)\sqrt{r}c\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \texttt{Pr}\bigg{[}\frac{(1-\Delta)}{(1+\Delta)}\sqrt{r}c\leq\big{\|}[\nabla_{W^{t}}\mathcal{L}_{i}^{t}]_{l,:}\big{\|}\leq\sqrt{r}c\bigg{]}\geq 1-p.$ (23)
Assuming a non-zero norm for all rows of $\nabla_{B^{t}}\mathcal{L}_{i}^{t}$ (for a sample $i$), we can always scale its rows so that the left condition is satisfied for all its rows. Then, the result on the right holds for the same sample $i$ and all its rows separately. If we do the same row scaling for all samples in a batch of size $b$, the right bound holds for all the samples simultaneously with probability at least $(1-nbp)$:
$\big{\|}[\nabla_{B^{t}}\mathcal{L}_{i}^{t}]_{l,:}\big{\|}=(1-\Delta)\sqrt{r}c\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (\forall\leavevmode\nobreak\ l\in[1,n],\forall i\in\mathcal{B}^{t})$
[…]
In other words, if we scale all the rows of sample gradients $\nabla_{B^{t}}\mathcal{L}_{i}^{t}$ in a batch to have norm $(1-\Delta)\sqrt{r}c$, then with probability at least $1-nbp$, all the sample gradients $\nabla_{W^{t}}\mathcal{L}_{i}^{t}$ in a batch have bounded frobenious norm $\sqrt{nr}c$. On the other hand, according to lemma 3.1, low rank adaptation adds a random noise to each row of $\nabla_{W^{t}}\mathcal{L}_{i}^{t}$ based on the norm of the row. More precisely, according to eq. 23, low-rank adaptation adds a Gaussian-like noise with variance at least $\frac{(\frac{(1-\Delta)}{(1+\Delta)}\sqrt{r}c)^{2}}{r}=\frac{(1-\Delta)^{2}}{(1+\Delta)^{2}}c^{2}$ to each element of the clipped sample gradient $\nabla_{W^{t}}\mathcal{L}_{i}^{t}$, whose frobenious norm was bounded in section 5. Also, according to lemma 4.3, this noise follows Gaussian distribution $\mathcal{N}(0,\frac{(1-\Delta)^{2}}{(1+\Delta)^{2}}c^{2})$ with probability $w_{g}$, where $(1-w_{g})\in\mathcal{O}\big{(}\frac{1}{\sqrt{mr}}\big{)}$.
5.1 Connecting LoRA/FLoRA to DPSGD Algorithm
As described above, when augmented with gradient scaling, the dynamics of LoRA/FLoRA is very close to DPSGD. However, it is not exactly the same: first, the distribution that the injected noise is sampled from is not exactly the pure Gaussian $\mathcal{N}(0,\frac{(1-\Delta)^{2}}{(1+\Delta)^{2}}c^{2})$. Also, as seen in section 5, the gradient scaling is probabilistic, while in DPSGD, the sample gradient clipping is deterministic, as if the upper-bound in section 5 always holds. Despite this, we can think of an intuitive relation to DPSGD. If we assume that the noise distribution is very close to Gaussian distribution (i.e., $w_{g}\approx 1$), and also $nbp\ll 1$, then we can consider the following interpretation of the low-rank adaptation of LoRA/FLoRA:
When scaling all the rows of sample gradients $\nabla_{B^{t}}\mathcal{L}_{i}^{t}$ to have norm $(1-\Delta)\sqrt{r}c$, low-rank adaptation adds a Gaussian noise with variance at least $\frac{(1-\Delta)^{2}}{(1+\Delta)^{2}}c^{2}$ to each element of the clipped sample gradients $\nabla_{W^{t}}\mathcal{L}_{i}^{t}$, whose frobenious norm is bounded by $\sqrt{nr}c$. This is equivalent to having a noise scale $z\geq\sqrt{b\frac{(1-\Delta)^{2}}{(1+\Delta)^{2}}c^{2}}/\sqrt{nr}c=\frac{(1-\Delta)}{(1+\Delta)}\sqrt{\frac{b}{nr}}$ for each batch of size $b$. The DP privacy parameters $\epsilon$ and $\delta$ resulting from this noise scale, which can be found by using a privacy accountant, e.g., moments accountant (Abadi et al., 2016), depend on the used batch size ratio (ratio of the batch size $b$ and the fine-tuning dataset size) and the number of steps $T$ taken during fine-tuning.
The connection drawn above is an approximate, yet meaningful, connection between LoRA/FLoRA and DPSGD, which provides a clear interpretation of what low-rank adaptation does. In fact, low-rank adaptation secretly mimics the mechanism of DPSGD approximately during fine-tuning. Hence, we expect it to provide robustness against privacy attacks against the fine-tuning data. Indeed, such a behavior has been observed implicitly in (Liu et al., 2024) for low-rank adaptation.
6 Conclusion
In this study, we establish an implicit connection between low-rank adaptation and differential privacy. We show that low-rank adaptation can be viewed as introducing random noise into the gradients w.r.t adapters coming from their fine-tuning with SGD. By quantifying the variance of this noise and bounding its deviation from pure Gaussian noise with the same variance, we demonstrate that low-rank adaptation, when combined with gradient clipping, approximates fine-tuning adapters with differential privacy. Although our theoretical analysis suggests that low-rank adaptation can provide implicit privacy similar to those of fine-tuning with differential privacy at a lower computational cost, empirical evaluation is necessary to fully validate these claims. In our ongoing future work, we will explore whether low-rank adaptation can effectively balance data privacy, security, and fine-tuning efficiency. Specifically, we aim to assess the practical performance of low-rank adaptation against security threats such as membership inference attacks (Zarifzadeh et al., 2024; Ye et al., 2022) and data extraction attacks (Carlini et al., 2019).
Acknowledgement
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Facebook award, and Google award. We also would like to thank Gautam Kamath (Cheriton School of Computer Science, University of Waterloo) and Abbas Mehrabian (Google DeepMind) for their valuable comments on some parts of this draft.
Appendix for on the Implicit Relation between Low-Rank Adaptation and Differential Privacy
Appendix A Useful Theorems
In this section, we mention some theorems, which we will use in our proofs.
Theorem A.1 (Chi-Squared distribution: [Mood and Franklin, 1974], Section 4.3, Theorem 7).
If the random variables $X_{i}$, $i=1,\ldots,k$, are normally and independently distributed with means $\mu_{i}$ and variances $\sigma_{i}^{2}$, then
$U=\sum_{i=1}^{k}\big{(}\frac{X_{i}-\mu_{i}}{\sigma_{i}}\big{)}^{2}$ (25)
has a chi-squared distribution with $k$ degrees of freedom: $U\sim\mathcal{X}_{k}^{2}$. Also, $\mathbb{E}[U]=k$ and $\texttt{Var}[U]=2k$.
The theorem above states that sum of the squares of $k$ standard normal random variables is a chi-squared distribution with $k$ degrees of freedom.
Lemma A.2 (Raw moment of Chi-Squared distribution).
Suppose $X\sim\mathcal{X}_{k}^{2}$. Then, the $m$-th raw moment of $X$ can be found as follows;
$\mathbb{E}[X^{m}]=\prod_{i=0}^{m-1}(k+2i)$ (26)
Proof.
From the definition of Chi-Squared distribution with $r$ degrees of reddom, $U$ has the following probability density function:
$f_{X}(x)=\frac{1}{2^{\frac{k}{2}}\Gamma(\frac{k}{2})}x^{\frac{k}{2}-1}e^{-\frac{x}{2}}$ (27)
Therefore, we have:
[…]
$=\prod_{i=0}^{m-1}(k+2i).$ (28)
Note that the fifth equality directly results from the property of gamma function that for $z>0$, $\Gamma(1+z)=z\Gamma(z)$. ∎
Theorem A.3 (Classical Central Limit Theorem: [Billingsley, 1995], Theorem 27.1).
Suppose that $\{X_{i}\}_{i=1}^{n}$, is an independent sequence of random variables having the same distribution with mean $\mu$ and positive variance $\sigma^{2}$. Define $S_{n}=\sum_{i=1}^{n}X_{i}$ as their sum. Let $Z_{n}$ be defined by
$Z_{n}=\frac{S_{n}-n\mu}{\sqrt{n}\sigma}.$ (29)
Then, the distribution of $Z_{n}$ approaches standard normal distribution as $n$ approaches infinity.
The theorem above states that $S_{n}$ is approximately, or asymptotically, distributed as a normal distribution with mean $n\mu$ and variance $n\sigma^{2}$.
The next theorem is about the Lindeberg’s condition, which is a sufficient (and under certain conditions also a necessary condition) for the Central Limit Theorem (CLT) to hold for a sequence of independent random variables $\{X_{i}\}_{i=1}^{n}$. Unlike the classical CLT stated above, which requires the sequence of random variables to have a finite variance and be both independent and identically distributed (i.i.d), Lindeberg’s CLT only requires the sequence of random variables to have finite variance, be independent and also satisfy the Lindeberg’s condition. The following states the theorem.
Theorem A.4 (Lindeberg and Lyapounov Theorem: [Billingsley, 1995], Theorem 27.2).
Suppose $X_{1},\ldots,X_{n}$ are $n$ independent random variables with $\mathbb{E}[X_{i}]=\mu_{i}$ and $\texttt{Var}[X_{i}]=\sigma_{i}^{2}>0$. Define $S_{n}=\sum_{i=1}^{n}X_{i}$ and let $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$. Also assume the following condition holds for all $\epsilon>0$:
[…]
where $P_{X_{i}}$ is the pdf of variable $X_{i}$. Assuming $Z_{n}=\frac{S_{n}-\sum_{i=1}^{n}\mu_{i}}{s_{n}}$, the distribution of $Z_{n}$ approaches standard normal distribution as $n$ approaches infinity.
The theorem above states that, given that Lindeberg’s condition is satisfied, $S_{n}$ is approximately, or asymptotically, distributed as a normal distribution with mean $\sum_{i=1}^{n}\mu_{i}$ and variance $s_{n}^{2}$, even if the sequence of variables are not identically distributed.
Appendix B Proofs
Using the theorems above, we are now able to prove lemma 3.1. See 3.1
Proof.
From the theorem’s assumption, we know that elements of $A$ are from $\mathcal{N}(0,\frac{1}{r})$. Therefore, we can rewrite the product $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ as the following product:
$q\cdot(\frac{A^{\top}A}{r}-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ (31)
where the elements of $A$ are now from standard normal distribution. Let $a_{i,j}$ denote the element in $i$-th row and $j$-th column of this new $A$. Therefore, for all $i$ and $j$, $a_{i,j}$ has distribution $\mathcal{N}(0,1)$. Let $B=\frac{A^{\top}A}{r}-\mathbb{I}_{m}$. Also, let $A_{i,:}$ and $A_{:,j}$ denote the $i$-th row and $j$-th column of the new $A$, respectively. We have:
$B_{i,i}=\frac{1}{r}[A^{\top}A]_{i,i}-1=\frac{1}{r}A_{:,i}^{\top}A_{:,i}-1=\frac{1}{r}\|A_{:,i}\|_{2}^{2}-1=(\frac{1}{r}\sum_{l=1}^{r}a_{l,i}^{2})-1$ (32)
From eq. 31, we know that $a_{l,i}$ is from standard normal distribution. Hence, $a_{l,i}^{2}$ is a chi-squared with 1 degree of freedom: $a_{l,i}^{2}\sim\mathcal{X}_{1}^{2}$. Therefore, $\sum_{l=1}^{r}a_{l,i}^{2}$, which is the sum of $r$ independent chi-squared variables with 1 degree of freedom, is a chi-squared with $r$ degrees of freedom: $\sum_{l=1}^{r}a_{l,i}^{2}\sim\mathcal{X}_{r}^{2}$ (see theorem A.1). Therefore, for $i\in\{1,\ldots,m\}$, we have:
$\mathbb{E}[B_{i,i}]=\mathbb{E}\big{[}\frac{\sum_{l=1}^{r}a_{l,i}^{2}}{r}\big{]}-1=\frac{r}{r}-1=0,$
$\texttt{Var}[B_{i,i}]=\texttt{Var}[\frac{\sum_{l=1}^{r}a_{l,i}^{2}}{r}]=\frac{\texttt{Var}(\mathcal{X}_{r}^{2})}{r^{2}}=\frac{2r}{r^{2}}=\frac{2}{r}.$ (33)
Similarly, we find the mean and variance of the non-diagonal elements $B_{i,j}(i\neq j)$ of $B$. We have:
$B_{i,j}=\frac{1}{r}[A^{\top}A]_{i,j}=\frac{1}{r}A_{:,i}^{\top}A_{:,j}=\frac{1}{r}\sum_{l=1}^{r}a_{l,i}a_{l,j},$ (34)
where $a_{l,i}$ and $a_{l,j}$ are independent and standard normal. Therefore, $a_{l,i}+a_{l,j}\sim\mathcal{N}(0,2)$. Similarly, $a_{l,i}-a_{l,j}\sim\mathcal{N}(0,2)$. So we can rewrite $a_{l,i}a_{l,j}$ as:
$a_{l,i}a_{l,j}=\frac{1}{4}(a_{l,i}+a_{l,j})^{2}-\frac{1}{4}(a_{l,i}-a_{l,j})^{2}=\frac{1}{2}z_{1}^{2}-\frac{1}{2}z_{2}^{2},$ (35)
where $z_{1}$ and $z_{2}$ are from standard normal. Therefore, $a_{l,i}a_{l,j}=\frac{\nu_{1}-\nu_{2}}{2}$, where $\nu_{1},\nu_{2}\sim\mathcal{X}_{1}^{2}$. Also, $a_{l,i}+a_{l,j}$ and $a_{l,i}-a_{l,j}$ are independent variables. Hence, $z_{1}$ and $z_{2}$ are independent, and likewise $\nu_{1}$ and $\nu_{2}$ are independent. We conclude that:
$a_{l,i}a_{l,j}=\frac{1}{2}(\nu_{1}-\nu_{2}),$ (36)
where $\nu_{1},\nu_{2}\sim\mathcal{X}_{1}^{2}$, and are independent.
Now, lets assume $\nu_{1},\nu_{2}\sim\mathcal{X}_{k}^{2}$ (a more general case), and let $M_{\nu_{1}}(t)=\mathbb{E}[e^{t\nu_{1}}]$ be the moment generating function (MGF) of $\nu_{1}$. In this case, we know that $M_{\nu_{1}}(t)=M_{\nu_{2}}(t)=(1-2t)^{-\frac{k}{2}}$ (MGF of $\mathcal{X}_{k}^{2}$). Hence, $M_{\nu_{1}-\nu_{2}}(t)=M_{\nu_{1}}(t)\cdot M_{\nu_{2}}(-t)=(1-4t^{2})^{-\frac{k}{2}}=\big{(}\frac{\frac{1}{4}}{\frac{1}{4}-t^{2}}\big{)}^{\frac{k}{2}}$, which is the MGF of a symmetric about origin variance-gamma distribution with parameters $\lambda=\frac{k}{2},\alpha=\frac{1}{2},\beta=0,\mu=0,\gamma=\frac{1}{2}$. Therefore, when $\nu_{1},\nu_{2}\sim\mathcal{X}_{k}^{2}$, then $\nu_{1}-\nu_{2}$ has this distribution, which has mean $\mu+2\beta\lambda/\gamma^{2}=0$ and variance $2\lambda(1+2\beta^{2}/\gamma^{2})/\gamma^{2}=4k$.
In eq. 36, we had $k=1$, as we had $\nu_{1},\nu_{2}\sim\mathcal{X}_{1}^{2}$. Hence, based on the discussion above, we have:
$\mathbb{E}[a_{l,i}a_{l,j}]$ $=0$ (37)
$\texttt{Var}[a_{l,i}a_{l,j}]$ $=\frac{1}{4}\texttt{Var}[\nu_{1}-\nu_{2}]=\frac{4k}{4}=1$ (38)
Consequently, based on eq. 34 and from the results above, we can compute the mean and variance of the non-diagonal elements of $B$ ($i\neq j$):
$\mathbb{E}[B_{i,j}]=\mathbb{E}\big{[}\frac{\sum_{l=1}^{r}a_{l,i}a_{l,j}}{r}\big{]}=\frac{\sum_{l=1}^{r}\mathbb{E}[a_{l,i}a_{l,j}]}{r}=0,$
$\texttt{Var}[B_{i,j}]=\texttt{Var}[\frac{\sum_{l=1}^{r}a_{l,i}a_{l,j}}{r}]=\frac{\sum_{l=1}^{r}\texttt{Var}[a_{l,i}a_{l,j}]}{r^{2}}=\frac{r}{r^{2}}=\frac{1}{r}.$ (39)
So far, we have computed the mean and variance of each entry in $B=\frac{A^{\top}A}{r}-\mathbb{I}_{m}\in\mathbb{R}^{m\times m}$ in appendix B and appendix B. Now, for a given $q\in\mathbb{R}^{1\times m}$, we have:
$q\cdot B=\sum_{l=1}^{m}q_{l}B_{l,:},$ (40)
where $B_{l,:}$ is row $l$ of $B$. Let $u_{i}$ denote the $i$-th element of $q\cdot B$. Hence, for each element $u_{i}$ ($i\in\{1,\ldots,m\}$), we have:
$\mathbb{E}[u_{i}]$ $=\mathbb{E}\bigg{[}\sum_{l=1}^{m}q_{l}B_{l,i}\bigg{]}=\sum_{l=1}^{m}q_{l}\mathbb{E}[B_{l,i}]=0,$
$\texttt{Var}[u_{i}]$ $=\texttt{Var}\bigg{[}\sum_{l=1}^{m}q_{l}B_{l,i}\bigg{]}=\sum_{l=1}^{m}q_{l}^{2}\texttt{Var}[B_{l,i}]=q_{i}^{2}\texttt{Var}[B_{i,i}]+\sum_{l\neq i}q_{l}^{2}\texttt{Var}[B_{l,i}]$
$=q_{i}^{2}\frac{2}{r}+\sum_{l\neq i}q_{l}^{2}\frac{1}{r}=\frac{q_{i}^{2}}{r}+\sum_{l=1}^{m}q_{l}^{2}\frac{1}{r}=\frac{q_{i}^{2}+\sum_{l=1}^{m}q_{l}^{2}}{r}\approx\frac{\sum_{l=1}^{m}q_{l}^{2}}{r}=\frac{\|q\|_{2}^{2}}{r},$ (41)
where the approximation is indeed valid because $m$, which is the dimension of the input of the current layer (see fig. 1), is a large integer. Finally, according to eq. 40, each element $u_{i}$ of $qB$ is the sum of $m$ random variables, for which the Lindeberg’s condition is also satisfied: as $m\to\infty$, $s_{m}^{2}=\frac{\|q\|_{2}^{2}}{r}\to\infty$ ($m$ is the dimension of $q$, and $s_{m}$ is the sum of variances of the $m$ random variables, which we found in appendix B). Hence, $[|u_{i}-0|>\epsilon s_{m}]\downarrow\emptyset$ as $m\to\infty$. Therefore, from theorem A.4, we also conclude that as $m\to\infty$, each element of $qB$ approaches a Gaussian with the mean and variance found in appendix B. Therefore, we conclude that having an $A$, where the elements of $A$ are i.i.d and from $\mathcal{N}(0,\frac{1}{r})$, then as $m\to\infty$, $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ approaches a Gaussian $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$, which completes the proof.
See 4.2
Proof.
From eq. 40, we had:
$u_{i}=\sum_{l\neq i,l=1}^{m}q_{l}B_{l,i}+q_{i}B_{i,i},$ (42)
where $B_{l,i}=\frac{1}{r}A^{\top}_{:,l}A_{:,i}=\frac{1}{2r}\sum_{t=1}^{r}V_{t}$, where $V_{t}\sim\texttt{Variance-Gamma}(\nu,\alpha,\beta,\mu)$ with $\nu=\beta=\mu=0$ and $\alpha=\frac{1}{2}$. Also $B_{i,i}=\frac{1}{r}A^{\top}_{:,i}A_{:,i}-1=\frac{\texttt{X}}{r}-1$, where $X\sim\mathcal{X}_{r}^{2}$. Therefore, we can rewrite the equation above for $u_{i}$ as:
$u_{i}=\sum_{l\neq i,l=1}^{m}\frac{q_{l}}{2r}\sum_{t=1}^{r}V_{t}+q_{i}(\frac{\texttt{X}}{r}-1)=\sum_{l\neq i,l=1}^{m}\sum_{t=1}^{r}\frac{q_{l}}{2r}V_{t}+\frac{q_{i}}{r}(\texttt{X}-r),$ (43)
where $V_{t}\sim\texttt{Variance-Gamma}(\nu,\alpha,\beta,\mu)$ with $\nu=\beta=\mu=0$ and $\alpha=\frac{1}{2}$ and $X\sim\mathcal{X}_{r}^{2}$. Hence, $V_{t}$ has mean 00 and variance $4$ and $(\texttt{X}-r)$ has mean 00 and variance $2r$. Also note that $X$ can be written as the summation of $r$ independent variables with distribution $\mathcal{X}_{1}^{2}$. Therefore, $u_{i}$ is the weighted sum of $mr$ independent random variables with mean 00. Also, from appendix B in the proof of lemma 3.1, we know that $u_{i}$ has mean 00 and variance $\frac{\|q\|_{2}^{2}}{r}$. Now, in order to bound the TV distance between the distribution of $u_{i}$ and $\mathcal{N}(0,\frac{\|q\|_{2}^{2}}{r})$, we have to use theorem 4.1 and eq. 14. More specifically, we have to find the third Lyapounov ratio $L_{3}=\frac{\sum_{i}\mathbb{E}[|X_{i}|^{3}]}{s_{n}^{3}}=\frac{\sum_{i}\mathbb{E}[|X_{i}|^{3}]}{\big{(}\sum_{i}\texttt{Var}[X_{i}]\big{)}^{3}}=\frac{\sum_{i}\mathbb{E}[|X_{i}|^{3}]}{\big{(}\sum_{i}\mathbb{E}[X_{i}^{2}]\big{)}^{3}}$, where $X_{i}$ is each of the $1+(m-1)r$ summands in eq. 43. First we note that, based on appendix B, $s_{n}^{3}=(\frac{\|q\|_{2}^{2}}{r})^{\frac{3}{2}}=\frac{\|q\|_{2}^{3}}{r\sqrt{r}}$. Now, we find the numerator $\sum_{i}\mathbb{E}[|X_{i}|^{3}]$. From [Gaunt, 2024], we know that for $V_{t}\sim\texttt{Variance-Gamma}(\nu,\alpha,0,0),\mathbb{E}[|V_{t}|^{r}]=\frac{2^{r}}{\sqrt{\pi}\alpha^{r}}\frac{\Gamma(\nu+(r+1)/2)\Gamma((r+1)/2)}{\Gamma(\nu+1/2)}$. Therefore, for $V_{t}\sim\texttt{Variance-Gamma}(0,\frac{1}{2},0,0),\mathbb{E}[|V_{t}|^{3}]=\frac{2^{6}}{\pi}$. On the other hand, we know that the skewness of $X\sim\mathcal{X}_{r}^{2}$ is equal to $\frac{\mathbb{E}[(X-\mathbb{E}[X])^{3}]}{\texttt{Var}[X]^{\frac{3}{2}}}=\frac{\mathbb{E}[(X-r)^{3}]}{(2r)^{\frac{3}{2}}}=\sqrt{\frac{8}{r}}$. Hence, $\mathbb{E}[(X-r)^{3}]=(2r)^{\frac{3}{2}}\sqrt{\frac{8}{r}}=8r$. Hence for $X\sim\mathcal{X}_{r}^{2}$, $\mathbb{E}[|X-r|^{3}]\geq\mathbb{E}[(X-r)^{3}]=8r$. Now, we can find the numerator $\sum_{i}\mathbb{E}[|X_{i}|^{3}]$ as:
$\sum_{i}\mathbb{E}[|X_{i}|^{3}]$ $=\sum_{l\neq i,l=1}^{m}\sum_{t=1}^{r}\frac{|q_{l}|^{3}}{8r^{3}}\mathbb{E}[|V_{t}|^{3}]+\frac{|q_{i}|^{3}}{r^{3}}\mathbb{E}[|\texttt{X}-r|^{3}]$
$=\sum_{l\neq i,l=1}^{m}\frac{|q_{l}|^{3}}{8r^{2}}\cdot\frac{2^{6}}{\pi}+\frac{|q_{i}|^{3}}{r^{3}}\mathbb{E}[|\texttt{X}-r|^{3}]$
$\approx\sum_{l\neq i,l=1}^{m}\frac{8|q_{l}|^{3}}{\pi r^{2}}+\frac{8|q_{i}|^{3}}{r^{2}}\approx\sum_{l=1}^{m}\frac{8|q_{l}|^{3}}{\pi r^{2}}=\frac{8}{\pi r^{2}}\|q\|_{3}^{3}.$ (44)
Therefore, for the sum $u_{i}$ in eq. 43, we have the third Lyapounov ratio:
$L_{3}=\frac{8}{\pi r^{2}}\|q\|_{3}^{3}\times\frac{r\sqrt{r}}{\|q\|_{2}^{3}}=\frac{8}{\pi\sqrt{r}}\bigg{(}\frac{\|q\|_{3}}{\|q\|_{2}}\bigg{)}^{3}.$ (45)
Therefore, based on theorem 4.1, we have:
$\|Q_{m}(x)-\Phi(x)\|_{TV}\leq\frac{8C_{D}}{\pi\sqrt{r}}\bigg{(}\frac{\|q\|_{3}}{\|q\|_{2}}\bigg{)}^{3},$ (46)
where $C_{D}\leq\frac{\pi\sqrt{r}}{8}$ is a constant, which depends only on $D$, where $D$ is an upperbound for the KL divergence between each of the random variable summands in eq. 43 and a Gaussian with the same mean and variance. Now, assuming $0<c\leq|q_{i}|\leq C$ for the elements $q_{i}$ in $q$, we have $\bigg{(}\frac{\|q\|_{3}}{\|q\|_{2}}\bigg{)}^{3}\leq\bigg{(}\frac{|C|}{|c|}\bigg{)}^{3}\frac{1}{\sqrt{m}}$. Therefore:
$\|Q_{m}(x)-\Phi(x)\|_{TV}\leq\frac{8C_{D}}{\pi}\bigg{(}\frac{|C|}{|c|}\bigg{)}^{3}\frac{1}{\sqrt{mr}}.$ (47)
Therefore,
$\|Q_{m}(x)-\Phi(x)\|_{TV}\in\mathcal{O}\bigg{(}\frac{1}{\sqrt{mr}}\bigg{)}.$ (48)