### v7

# Low-Rank Adaptation Secretly Imitates Differentially Private SGD

Abstract
As pre-trained language models grow in size, full fine-tuning their parameters on task adaptation data becomes increasingly impractical. To address this challenge, some methods for low-rank adaptation of language models have been proposed, e.g. LoRA, which incorporates trainable low-rank decomposition matrices into only some parameters of the pre-trained model, called adapters. This approach significantly reduces the number of trainable parameters compared to fine-tuning all parameters or adapters. In this work, we look at low-rank adaptation method from the lens of data privacy. We show theoretically that the low-rank adaptation used in LoRA is equivalent to fine-tuning adapters with noisy batch gradients - just like what DPSGD algorithm does. We also quantify the variance of the injected noise as a decreasing function of adaptation rank. By establishing a Berry-Esseen type bound on the total variation distance between the injected noise distribution and a Gaussian noise distribution with the same variance, we show that the dynamics of low-rank adaptation is very close to when DPSGD is performed w.r.t the adapters. Following our theoretical findings and approved by our experimental results, we show that low-rank adaptation provides robustness to membership inference attacks w.r.t the fine-tuning data.
1 Introduction
Stochastic Gradient Descent (SGD) is the power engine of training deep neural networks, which updates parameters of a model by using batch gradients. Modern deep learning models, e.g. GPT-3 (gpt3, ) and Stable Diffusion (stable_diffusion, ), have a large number of parameters, which induces a large space complexity for their fine-tuning with SGD. Using more advanced methods like momentum technique (momentum, ) and gradient accumulation (gradacc, ) exacerbates this space complexity (JMLR:v12:duchi11a, ). Some prior works tried to reduce the space complexity of fine-tuning large models by tuning a subset of parameters and storing the information about only a portion of them (houlsby, ; zaken, ). LoRA (lora, ) is such an algorithm, which only updates some of the parameter, called adapters, by restricting their updates to be a low-rank matrix. This low-rank restriction considerably reduces the number of trainable parameters, at the cost of limiting the optimization space of the adapters. As another parameter-efficient training technique, ReLoRA (relora, ) utilizes low-rank updates to train high-rank networks to mitigate the low-rank constraint of LoRA. Similarly, the work in (flora, ) identifies that the dynamics of LoRA can be approximated by a random matrix projection. Based on this, the work proposes FLoRA to achieve high-rank updates by resampling the random projection matrices, which enables achieving higher utility while still enjoying the sublinear space complexity of LoRA.
On the other hand, from the lens of data privacy, fine-tuning data often happens to be privacy-sensitive. In such scenarios, differentially private (DP) fine-tuning algorithms are usually used to provide rigorous privacy guarantees w.r.t the data. DP full fine-tuning runs DPSGD (abadi, ) to fine-tune the entire existing parameters of a pre-trained model. However, due to computing sample gradients and clipping them for every data sample in a batch, running DPSGD on all parameters induces high space complexities, even higher than non-private fine-tuning of all parameters. As a remedy, to achieve both data privacy and low space complexity, DPSGD can be applied to only the adapters in the pre-trained model at fine-tuning time. That applying DPSGD to the adapters can provide privacy w.r.t the fine-tuning data can be mostly attributed to its random noise injection into batch gradients.
In this work, we draw a connection between low-rank adaptation and DPSGD w.r.t the adapters. We show that low-rank adaptation of the adapters is equivalent to fine-tuning them with noisy batch gradients, which is very close to what DPSGD does for fine-tuning adapters privately. We also quantify the variance of the injected noise as a function of adaptation rank and show that it increases as the adaptation rank decreases. Furthermore, we show that the distribution of the injected noise is very close to a Gaussian distribution with the same variance. Our derivations show that low-rank adaptation and DPSGD w.r.t the adapters are very close to each other in terms of their mechanisms. This implies that low rank adaptation can potentially provide an informal privacy w.r.t the fine-tuning data without inducing the high space complexity of running DPSGD on adapters or all parameters. This behavior from LoRA was implicitly observed in a prior work (liu2024precuriousinnocentpretrainedlanguage, ) without a clear understanding of its root reason. In this work, we shed light on the reason behind this behavior of low-rank adaptation with theoretical and experimental results. The highlights of our contributions are as follows:
  • We show that low-rank adaptation of adapter parameters is equivalent to fine-tuning them with noisy batch gradients (eq. 10).
  • We quantify the variance of the noise injected into the batch gradients w.r,t the adapters, and show that the injected noise approaches a Gaussian distribution as the input dimension of the adaptation layer increases (lemma 3.1).
  • We bound the total variation distance between the distributions of the injected noise and a pure Gaussian noise with the same variance. The bound decreases as the input dimension of the adaptation layer and the adaptation rank increase (lemma 4.2).
  • Baed on the above findings, we show that the dynamics of low-rank adaptation is very close to that of fine-tuning the adapters with DPSGD (Section 5).
  • Finally, we show experimentally that low-rank adaptation provides robustness to memebrship inference atatcks w.r.t the fine-tuning dataset (Section 6), confirming our theoretical findings.
2 Dynamics of low-rank task adaptation
We start by studying the dynamics of low-rank adaptation. In order to update a pre-trained adapter weight $W\in\mathbb{R}^{n\times m}$ located in a layer of a pre-trained model, LoRA incorporates low-rank decomposition matrices $B\in\mathbb{R}^{n\times r}$ and $A\in\mathbb{R}^{r\times m}$, where $r\ll\min\{n,m\}$, and performs the following forward pass in an adaptation layer (see fig. 1):
$y=(W+BA)x=Wx+BAx,$ (1)
Figure 1: Low-rank decomposition of LoRA in an adaptation layer.
where $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{n}$ are the input and the pre-activation output of the current adaptation layer, respectively. For instance, there are 12 self-attention layers in GPT-2 model and when fine-tuning it with LoRA, $W$ can be considered to be the query layer in each of those 12 layers, which has dimension $(m,n)=(768,2304)$. For each adaptation layer, it is common to initialize $B$ with an all-zero matrix and sample $A$ from the normal distribution $\mathcal{N}(0,\frac{1}{r})$. We will use this intialization of $A$ for deriving our theoretical results. When back-propagating, LoRA calculates the gradients w.r.t only $A$ and $B$, which can be found as follows:
$\frac{\partial\mathcal{L}}{\partial A}$ $=\frac{\partial BA}{\partial A}\cdot\frac{\partial\mathcal{L}}{\partial BA}=B^{\top}\cdot\frac{\partial\mathcal{L}}{\partial y}\cdot\frac{\partial y}{\partial BA}=B^{\top}\cdot\frac{\partial\mathcal{L}}{\partial y}\cdot x^{\top}=B^{\top}(\nabla_{W}\mathcal{L}),$
$\frac{\partial\mathcal{L}}{\partial B}$ $=\frac{\partial\mathcal{L}}{\partial BA}\cdot\frac{\partial BA}{\partial B}=\frac{\partial\mathcal{L}}{\partial y}\cdot\frac{\partial y}{\partial BA}\cdot A^{\top}=\frac{\partial\mathcal{L}}{\partial y}\cdot x^{\top}\cdot A^{\top}=(\nabla_{W}\mathcal{L})A^{\top}.$ (2)
Hence, $\frac{\partial\mathcal{L}}{\partial A}\in\mathbb{R}^{r\times m}$ and $\frac{\partial\mathcal{L}}{\partial B}\in\mathbb{R}^{n\times r}$. In fact, LoRA down-projects the batch gradient $\nabla_{W}\mathcal{L}$ from $\mathbb{R}^{n\times m}$ to a lower dimension, and updates the matrices $A$ and $B$ with the resulting projections of $\nabla_{W}\mathcal{L}$. Let $A^{0}$, $B^{0}$ and $W^{0}$ denote the initial values of the matrices $A$, $B$ and the adapter $W$ (coming from the pretrained model). We restate the following thorem from (flora, ) without its proof:
Theorem 2.1 (flora ).
Let LoRA update matrices $A$ and $B$ with SGD as:
$A^{t+1}\leftarrow A^{t}-\eta\frac{\partial\mathcal{L}}{\partial A^{t}}=A^{t}-\eta B^{t}{{}^{\top}}(\nabla_{W^{t}}\mathcal{L}^{t}),$ (3)
$B^{t+1}\leftarrow B^{t}-\eta\frac{\partial\mathcal{L}}{\partial B^{t}}=B^{t}-\eta(\nabla_{W^{t}}\mathcal{L}^{t})A^{t}{{}^{\top}},$ (4)
where $\eta$ is the learning rate. We assume $\|\sum_{t=0}^{T}\nabla_{W^{t}}\mathcal{L}^{t}\|_{F}\leq L$ for every $T$, implying that the model stays within a finite Euclidean ball from $W^{0}$. In this case, the dynamics of $A^{t}$ and $B^{t}$ are given by
$A^{t}=A^{0}+\eta A^{0}f_{A}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ B^{t}=\eta f_{B}(t)A^{0}{{}^{\top}},$ (5)
where the forms of $f_{A}(t)\in\mathbb{R}^{m\times m}$ and $f_{B}(t)\in\mathbb{R}^{n\times m}$ are expressed in the proof. In particular, $\|f_{A}(t)\|_{2}\leq\frac{\eta L^{2}\big{(}1-(\eta^{2}L^{2})^{t}\big{)}}{1-\eta^{2}L^{2}}$ for every $t$.
In the equation above, $W^{t}=W^{0}+B^{t}A^{t}$, i.e. the equivalent forward pass parameter at time step $t$. Let us denote the total changes of $A$ and $B$ after $T$ gradient steps as $\Delta A$ and $\Delta B$, respectively. Then, the forward pass in eq. 1 will get updated to:
$\big{(}W^{0}+(B^{0}+\Delta B)(A^{0}+\Delta A)\big{)}x=\big{(}W^{0}+\Delta BA^{0}+\Delta B\Delta A\big{)}x,$ (6)
where we have substituted the initialization $B^{0}=\mathbf{0}\in\mathbb{R}^{n\times r}$. From eq. 5 and after substituting the values of $\Delta A$ and $\Delta B$ after $T$ gradient steps into the above equation, we get to the following updated forward pass parameter for the adaptation layer:
$W^{0}+\Delta BA^{0}+\Delta B\Delta A=W^{0}+\eta f_{B}(T)A^{0}{{}^{\top}}A^{0}+\eta^{2}f_{B}(T)A^{0}{{}^{\top}}A^{0}f_{A}(T).$ (7)
Also, from theorem 2.1, we have $\|f_{A}(T)\|_{2}\leq\|f_{A}(T)\|_{F}\leq\frac{\eta L^{2}\big{(}1-(\eta^{2}L^{2})^{T}\big{)}}{1-\eta^{2}L^{2}}$. Hence, if $\eta\ll 1/L$, we have $\lim_{T\to\infty}\eta\|f_{A}(T)\|_{2}=\lim_{T\to\infty}\frac{(\eta L)^{2}\big{(}1-(\eta L)^{(2T)}\big{)}}{1-(\eta L)^{2}}\ll 1$. Therefore, the last term on the right side is significantly smaller than the second term $\eta f_{B}(T)A^{0}{{}^{\top}}A^{0}$. Hence, the second term is the dominant term in the update to the forward pass parameter $W^{0}$. Therefore, we can closely approximate LoRA by freezing $A$ to its initial value $A^{0}$ and adapting only the matrix $B$. In this case:
$W^{0}+\Delta BA^{0}+\Delta B\Delta A$ $=W^{0}+\Delta BA^{0}=W^{0}+\eta\tilde{f}_{B}(T)A^{0}{{}^{\top}}A^{0},$ (8)
where $\tilde{f}_{B}(0)=\mathbf{0}$ and $\tilde{f}_{B}(t+1)=\tilde{f}_{B}(t)-\nabla_{W^{t}}\mathcal{L}^{t}$. Equivalently, $\tilde{f}_{B}(T)=-\sum_{t=0}^{T-1}\nabla_{W^{t}}\mathcal{L}^{t}$. Substituting this into the equation above, we get:
$W^{0}+\Delta BA^{0}+\Delta B\Delta A=W^{0}+\Delta BA^{0}=W^{0}-\eta\sum_{t=0}^{T-1}\big{[}(\nabla_{W^{t}}\mathcal{L}^{t})A^{0}{{}^{\top}}A^{0}\big{]},$ (9)
where the last term shows the exact change in the forward pass parameter $W^{0}$ after performing $T$ rounds of SGD on the matrix $B$. The term $\nabla_{W^{t}}\mathcal{L}^{t}$ in the sum is the batch gradient that would be obtained if we had “fine-tuned" the adapter $W$ using SGD. Therefore, low rank adaptation with LoRA can be viewed as performing a random projection of stochastic batch gradient $\nabla_{W^{t}}\mathcal{L}^{t}$ in every step $t$ with matrix $A^{0}{{}^{\top}}$ and projecting it back with matrix $A^{0}$.
Having understood the connection between low-rank adaptation in LoRA and random projection, in the next section, we show that the update in eq. 9 to the forward pass parameter $W$ is equivalent to “fine-tuning" adapter $W$ with a noisy version of batch gradient $\nabla_{W^{t}}\mathcal{L}^{t}$ at every time step $t$. This is our first step towards establishing a comparison framework between the dynamics of low-rank adaptation and that of DPSGD w.r.t the dapter $W$.
3 Random noise injection by low-rank adaptation
For simplicity of our analysis, we keep using the approximation of LoRA shown in the previous section, i.e. a fixed projection matrix $A^{0}$. As illustrated in eq. 10, the update in eq. 9 to the initial forward pass parameter $W^{0}$ after $T$ SGD steps w.r.t $B^{t}$ can be rewritten as:
$W^{0}+\Delta BA^{0}=W^{0}-\eta\sum_{t=0}^{T-1}\bigg{[}\underbrace{\nabla_{W^{t}}\mathcal{L}^{t}}_{\textit{ batch gradient w.r.t $W^{t}$}}+\underbrace{\nabla_{W^{t}}\mathcal{L}^{t}(A^{0}{{}^{\top}}A^{0}-\mathbb{I}_{m})}_{\textit{noise}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathbb{R}^{n\times m}}\bigg{]},$ (10)
The second term in the sum represents a noise term introduced by the low-rank adaptation. Therefore, “low-rank adaptation" of pretrained adapter $W^{0}$ by introducing low-rank matrices $A$ and $B$ is equivalent to “fine-tuning" it with noisy batch gradients. We are now particularly interested in the behavior of this noise term, which is added to each batch gradient $\nabla_{W^{t}}\mathcal{L}^{t}$ in every step $t$. Recall that the entries of $A^{0}$ were sampled from $\mathcal{N}(0,\frac{1}{r})$ (see fig. 1), and that each of the $r$ columns of $A^{0}{{}^{\top}}$ is an $m$-dimensional Gaussian random variable. Consequently, $A^{0}{{}^{\top}}A^{0}$ follows a Wishart distribution with $r$ degrees of freedom (Wishart1, ), which is the multivariate generalization of the chi-squared distribution. Therefore, for any $q\in\mathbb{R}^{1\times m}$, $q\cdot(A^{0^{\top}}A^{0}-\mathbb{I}_{m})$ is a weighted sum of multiple chi-squared random variables, which implies that the result follows a Gaussian distribution approximately, according to the Central Limit Theorem (CLT) (bhattacharya2016course, ). Therefore, we first state and prove the following lemma concerning the noise term in eq. 10.
Lemma 3.1.
Let $A\in\mathbb{R}^{r\times m}$ be a matrix with i.i.d entries sampled from $\mathcal{N}(0,\frac{1}{r})$. Given a fixed $q\in\mathbb{R}^{1\times m}$, the distributions of the elements of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ approach the Gaussian distribution $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$, as $m$ increases.
The result above can be extended to matrices multiplication, as in eq. 10: for $A\in\mathbb{R}^{r\times m}$ sampled from $\mathcal{N}(0,\frac{1}{r})$ and $Q\in\mathbb{R}^{n\times m}$ and as $m$ grows, the product $G=Q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{n\times m}$ approaches a Gaussian distribution, where $G_{i,j}$ ($1\leq i\leq n$) has distribution $\mathcal{N}(0,\|[Q]_{i,:}\|^{2}/r)$, where $[Q]_{i,:}$ is the $i$-th row of $Q$.
The lemma above shows that the last term in eq. 10 can indeed be looked at as a random noise term, whose variance decreases with $r$. Although lemma 3.1 was proved when $m$ approaches infinity, in practical scenarios $m$ is limited (the input dimension of the adaptation layer). Hence, the distribution of the injected noise is not a pure Gaussian. In the next section, we bound the deviation of the injected noise distribution from a pure Gaussian distribution with the same variance.
4 Bounding the distance to the Gaussian distribution
Having proved lemma 3.1 when $m$ approaches infinity, yet we need to quantify the distance between the distribution of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ (for a given $q\in\mathbb{R}^{1\times m}$) and the Gaussian distribution with the same variance for limited values of $m$. In this section, we derive a closed form upper-bound for the total variation distance between the distribution of each element of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ and the Gaussian distribution $\mathcal{N}(0,\|q\|^{2}/r)$.
Suppose $X_{1},\ldots,X_{n}$ are $n$ independent random variables that are not necessarily i.i.d, and $\mathbb{E}[X_{i}]=0$ (mean) and $\texttt{Var}[X_{i}]=\sigma_{i}^{2}>0$ (variance). Define $S_{n}=\sum_{i=1}^{n}X_{i}$ and let $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$. Assuming $Z_{n}=S_{n}/s_{n}$, and having Lindeberg’s condition satisfied (see theorem B.3 and theorem B.4 in the appendix), the normalized sum $Z_{n}$ has standard normal distribution in a weak sense for a bounded $n$. More precisely, the closeness of the cumulative distribution function (CDF) $F_{n}(x)=\texttt{Pr}\{Z_{n}\leq x\}$ to the standard normal CDF has been studied in terms of the Lyapounov ratios:
$L_{t}=\frac{\sum_{i=1}^{n}\mathbb{E}[|X_{i}|^{t}]}{s_{n}^{t}}.$ (11)
Particularly, if all $\{X_{i}\}_{i=1}^{n}$ are i.i.d and have a finite third absolute moment $\mathbb{E}[|X_{i}|^{3}]$, the classical Berry-Esseen theorem (Esseen1945FourierAO, ; feller1971, ; Petrov1975, ) bounds the Kolmogrov distance between $F_{n}(x)$ and $\Phi(x)$:
$\sup_{x}|F_{n}(x)-\Phi(x)|\leq CL_{3},$ (12)
where $C$ is an absolute constant. In the more general case, when $\{X_{i}\}_{i=1}^{n}$ are independent (and not necessarily i.i.d), the number of summand variables $n$ implicitly affects the value of $L_{3}$. Yet, we can bound the difference between $F_{n}(x)$ and $\Phi(x)$ in terms of generally stronger distances of total variation or entropic distance (Bobkov2011, ). To this end, let $D(X_{i})$ denote the KL divergence between $X_{i}$ and a normal variable from $\mathcal{N}(0,\sigma_{i}^{2})$, i.e. the KL divergence between $X_{i}$ and a normal variable with the same variance. Therefore, we have the following theorem about the distance between $F_{n}$ and $\Phi$:
Theorem 4.1 ((Bobkov2011, ), theorem 1.1).
Assume that the independent random variables $X_{1},\ldots,X_{n}$ have finite third absolute moments, and that $D(X_{i})\leq D$, where $D$ is a non-negative number. Then,
$\|F_{n}(x)-\Phi(x)\|_{TV}\leq C_{D}L_{3},$ (13)
[…]
Having the theorem above, we can now derive a Berry-Esseen type bound for the total variation distance between each element of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ in lemma 3.1 and the normal law $\mathcal{N}(0,\|q\|^{2}/r)$: we need to find the third Lyapounov ratio for the summands contributing to each of the $m$ elements in $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$. To this end, we state and prove the following lemma:
Lemma 4.2.
Let $A\in\mathbb{R}^{r\times m}$ be a matrix with i.i.d entries sampled from $\mathcal{N}(0,\frac{1}{r})$. Given a fixed $q\in\mathbb{R}^{1\times m}$ with $0<c\leq|q_{i}|\leq C$, let $u=q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$. Let $u_{i}$ be the $i$-th element of $u$ and $Q_{m}(x)=\texttt{Pr}\{u_{i}\leq x\}$, the CDF of $u_{i}$. Also, let $\Phi(x)$ be the CDF of $z\sim\mathcal{N}(0,\frac{\|q\|^{2}}{r})$. Then:
$\|Q_{m}(x)-\Phi(x)\|_{TV}\in\mathcal{O}\bigg{(}\frac{1}{\sqrt{mr}}\bigg{)}.$ (14)
The above result shows that the distribution of $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ gets closer and closer to $\mathcal{N}(0,\|q\|^{2}/r)$ as $m$ and $r$ increase. We can now benefit from the useful coupling characterization of the total variation distance to establish a more understandable interpretation of the above result.
The coupling characterization of the total variation distance.
For two distributions $P$ and $Q$, a pair of random variables $(X,Y)$, which are defined on the same probability space, is called a coupling for $P$ and $Q$ if $X\sim P$ and $Y\sim Q$ (Levin2008MarkovCA, ; abbas, ). A very useful property of total variation distance is the coupling characterization (Levin2008MarkovCA, ):
$\|P-Q\|_{TV}\leq t$ if and only if there exists a coupling $(X,Y)$ for them such that $\texttt{Pr}\{X\neq Y\}\leq t$.
We can use the coupling characterization to prove the following lemma from lemma 4.2.
Lemma 4.3.
Let $A\in\mathbb{R}^{r\times m}$ be a matrix with i.i.d entries sampled from $\mathcal{N}(0,\frac{1}{r})$. Given a fixed $q\in\mathbb{R}^{1\times m}$ with $0<c\leq|q_{i}|\leq C$, let $u=q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$. Let $u_{i}$ be the $i$-th element of $u$. There exists a coupling $(u_{i},z)$, where $z\sim\mathcal{N}(0,\frac{\|q\|^{2}}{r})$ and
$\texttt{Pr}\{u_{i}\neq z\}\in\mathcal{O}\bigg{(}\frac{1}{\sqrt{mr}}\bigg{)}.$ (15)
The lemma above means that each element of $q\cdot(A^{\top}A-\mathbb{I}_{m})$ follows a mixture of distributions: $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$ with weight $w_{g}$ and another distribution $M$ which we dont know - with weight $(1-w_{g})\in\mathcal{O}\big{(}\frac{1}{\sqrt{mr}}\big{)}$. For instance, when finetuning GPT-2 model with LoRA with $r=16$, $1/\sqrt{mr}\approx 9\times 10^{-3}$, suggesting that the distribution of the injected noise is indeed very close to a Gaussian distribution. The larger $mr$, the closer the mixture distribution gets to pure Gaussian distribution $\mathcal{N}(0,\|q\|^{2}/r)$. Having established the results above, we can now make a clear comparison between the dynamics of low-rank adaptation of $W$ and the dynamics of DPSGD w.r.t $W$.
5 Comparing Low Rank Adaptation and DPSGD
According to section 2, the matrix $B^{0}$ after performing $T$ SGD steps w.r.t $B$ will be updated to:
$B^{0}\leftarrow B^{T}=-\eta\sum_{t=0}^{T-1}\nabla_{B^{t}}\mathcal{L}^{t}=-\eta\sum_{t=0}^{T-1}\nabla_{W^{t}}\mathcal{L}^{t}A^{0}{{}^{\top}}.$ (16)
In that case, according to eq. 9, the forward-pass parameter $W^{0}$ will be updated to:
$W^{0}\leftarrow W^{0}+B^{T}A^{0}=W^{0}-\eta\sum_{t=0}^{T-1}\nabla_{W^{t}}\mathcal{L}^{t}A^{0}{{}^{\top}}A^{0}.$ (17)
When performing LoRA, we never compute $\nabla_{W^{t}}\mathcal{L}^{t}$ due to its memory overhead, but rather we just compute $\nabla_{B^{t}}\mathcal{L}^{t}$ at each SGD step $t$ and compute the left side of eq. 16. However, as we showed earlier, this is mathematically equivalent to computing $\nabla_{W^{t}}\mathcal{L}^{t}$ at each step $t$ and computing the right side of eq. 17. Therefore, in this section, we can assume that LoRA computes $\nabla_{W^{t}}\mathcal{L}^{t}$ at each step $t$ and performs the update in eq. 17. As we learned earlier and based on our understandings from lemma 4.3, low rank adaptation (with rank $r$) of adapter $W\in\mathbb{R}^{n\times m}$ at time step $t$ is equivalent to fine-tuning it with the noisy stochastic batch gradients ~$\tilde{\nabla}_{W^{t}}\mathcal{L}^{t}=\nabla_{W^{t}}\mathcal{L}^{t}+N^{t}$, where $N^{t}\in\mathbb{R}^{n\times m}$ is a noise-term with Gaussian-like distribution: $\texttt{Pr}\{N_{i,j}^{t}\neq z_{i}^{t}\}\in\mathcal{O}\big{(}1/\sqrt{mr}\big{)}$, where $z_{i}^{t}\sim\mathcal{N}(0,\frac{\|[\nabla_{W^{t}}\mathcal{L}^{t}]_{i,:}\|^{2}}{r})$. In other words, low-rank adaptation adds noise to each row of batch gradient $\nabla_{W^{t}}\mathcal{L}^{t}$, and the standard deviation of the noise added to the elements of the row $i$ is proportional to the $\ell_{2}$ norm of row $i$. Also, as $mr$ grows, the distribution of noise element $N_{i,j}^{t}$ gets closer to $\mathcal{N}(0,\frac{\|[\nabla_{W^{t}}\mathcal{L}^{t}]_{i,:}\|^{2}}{r})$.
The operation above is very similar to what DPSGD (abadi, ) does when it is applied to $W$ (see appendix A for a short background about DPSGD). At the $t$-th gradient update step on a current adapter parameter $W^{t}$, DPSGD computes the following noisy clipped batch gradient on a batch $\mathcal{B}^{t}$ with size $b$:
~$\tilde{\nabla}_{W^{t}}\mathcal{L}^{t}=\bigg{(}\frac{1}{b}\sum_{i\in\mathcal{B}^{t}}\bar{\nabla}_{W^{t}}\mathcal{L}_{i}^{t}\bigg{)}+\mathcal{N}(0,\frac{(c^{t})^{2}z^{2}}{b^{2}}),$ (18)
where ¯$\bar{\nabla}_{W^{t}}\mathcal{L}_{i}^{t}$ is the clipped sample gradient computed on sample $i$ with clipping threshold $c^{t}$ at time step $t$. If $c^{t}\geq\|{\nabla}_{W^{t}}\mathcal{L}_{i}^{t}\|$, then the clipping will be ineffective and can be ignored. On the other hand, following eq. 17, after averaging over a batch of samples $\mathcal{B}^{t}$, LoRA gets to the following noisy batch gradient to update the forward pass parameter $W^{t}$ in the adaptation layer at time step $t$:
~$\tilde{\nabla}_{W^{t}}\mathcal{L}^{t}=\frac{1}{b}\bigg{[}\sum_{i\in\mathcal{B}^{t}}\nabla_{W^{t}}\mathcal{L}_{i}^{t}\leavevmode\nobreak\ A^{0}{{}^{\top}}A^{0}\bigg{]}=\bigg{(}\frac{1}{b}\sum_{i\in\mathcal{B}^{t}}\nabla_{W^{t}}\mathcal{L}_{i}^{t}\bigg{)}+\bigg{(}\frac{1}{b}\sum_{i\in\mathcal{B}^{t}}\nabla_{W^{t}}\mathcal{L}_{i}^{t}\bigg{)}\big{(}A^{0}{{}^{\top}}A^{0}-\mathbb{I}_{m}\big{)}.$ (19)
Very similar to eq. 18, the right side consists of a batch gradient term and a noise term. The main difference between the noise addition mechanism in low-rank adaptation (eq. 19) and that in DPSGD (eq. 18) is that DPSGD adds a pure Gaussian noise with variance $(c^{t})^{2}z^{2}/b^{2}$ to all elements of the average clipped batch gradient at time step $t$. In contrast, LoRA adds random noise to different rows of the average batch gradient according to their norms: rows with larger norms experience more noise. Also, the injected noise to each element follows a Gaussian distribution with a high probability.
The interpretation above is a meaningful comparison between the mechanisms of LoRA and DPSGD, which provides a clear interpretation of what low-rank adaptation does. In fact, low-rank adaptation secretly mimics the mechanism of DPSGD by using noisy batch gradients w.r.t the adapter parameters during the fine-tuning time. Accordingly, we expect it to provide informal robustness to privacy attacks against the fine-tuning data, just like DPSGD. Our theoretical findings suggest that the smaller the adaptation rank, the more robust the fine-tuned model is expected to be to privacy attacks.
6 Experiments
In this section, we perform membership inference attacks (MIA) on some fine-tuned models to show how low-rank adaptation indeed provides robustness against them.
Datasets: We run our experiments on three fine-tuning datasets with confidential properties, including Penn Treebank (PTB)(marcusetal-1993-building, ), Enron (Klimt2004TheEC, ) and Pubmed (cohan2018discourseawareattentionmodelabstractive, ). We split each dataset into three parts and use them as train, validation and auxilliary datasets.
[…]
Membership Inference Attacks: We use a calibrated membership score for membership inference (carlini2022membershipinferenceattacksprinciples, ; mireshghallah2022memorizationnlpfinetuningmethods, ; sablayrolles2019whiteboxvsblackboxbayes, ; ye2022enhancedmembershipinferenceattacks, ). More specifically, we train a reference model $\theta_{ref}$ on the auxiliary dataset and use the following signal for classifying a sample $x$ as a member/non-member of the fine-tuning dataset:
$I_{\theta}(x):=\mathbb{I}[\mathcal{L}(x;\theta)-\mathcal{L}(x;\theta_{ref})<\gamma].$ (20)
The threshold $\gamma$ is set to the highest value for which the false positive rate over all samples will be lower than $\alpha$. We consider $\alpha\in\{10\%,1\%\}$ and report the corresponding true positive rates (TPR).
Metrics: To measure utility, we use the perplexity on the validation set (PPL). For measuring the effectiveness of MIA, we use AUC and TPR@@@@FPR $\alpha$ for $\alpha\in\{10\%,1\%\}$.
Table 1: Membership inference attack on GPT-2. LoRA (with both frozen and trainable $A$) shows robustness to the attacks in terms of AUC and TPR of MIA by mimicking DPSGD and incorporating noisy batch gradients w.r.t adapters during adaptation.
Dataset Enron Pubmed PTB
Evaluation Criterion Utility MIA Success Metrics Utility MIA Success Metrics Utility MIA Success Metrics
[…]
Full fine-tuning 18.49 0.876 66.31 13.02 16.64 0.929 80.67 6.48 27.70 0.963 90.54 65.84
LoRA ($r=16$) 19.66 0.757 40.56 8.08 18.82 0.816 50.43 2.68 31.18 0.961 89.97 59.88
LoRA (frozen $A^{0}$, $r=16$) 21.90 0.614 16.76 4.49 20.85 0.546 14.93 2.37 37.93 0.952 82.23 55.58
DPSGD on adapters ($\epsilon=20$) 27.74 0.541 13.62 1.19 27.16 0.532 12.21 1.78 44.53 0.542 55.21 9.87
Experimental results: In Table 1, we have reported the results obtained from evaluating MIA effectiveness on the GPT-2 model and different fine-tuning algorithms. As observed, full fine-tuning all parameters achieves the best utility (PPL) on all datasets with the cost of a considerable vulnerability to MIA. In contrast, choosing a subset of parameters in the existing 12 layers of the model (the adapters) and adapting them with LoRA improves robustness to MIA considerably. As observed, LoRA achieves a considerably lower AUC and TPR@FPR compared to full fine-tuning at the cost of a slight drop in utility (larger perplexity), thanks to the low-rank adaptation mechanism incorporating noisy batch gradients w.r.t to the adapter parameters. Also, we can observe that LoRA with a frozen $A^{0}$ enhances this robustness even further. For comparison, we have included the results for fine-tuning the same adapter parameters in the existing 12 layers with DPSGD. As expected, fine-tuning with DPSGD also improves the robustness to MIA compared to full fine-tuning. However, it leads to a more severe drop in utility, which can be attributed to the fact that DPSGD adds noise uniformly with the same variance to all the elements in batch gradients. In contrast, LoRA tends to add less noise to the rows in batch gradients that have a smaller norm. It is also noteworthy that, in all our experiments, running DPSGD needed more than twice as much runtime as LoRA needed, due to its higher space complexity. Figure 2 summarizes the results in table 1 more clearly.
Figure 2: The trade-off existing between utility, privacy and computational/space complexity (run time) for different adaptation algorithms. LoRA provides robustness against privacy attacks to fine-tuning data without much adverse effect on utility or additional computational/space complexity.
Figure 3: The ffect of adaptation rank $r$ on the model fine-tuned on Pubmed averaged over three different data splits (random seeds). Left: The perplexity (utility) of the fine-tuned model decreases (increases) as $r$ increases. Middle: The AUC of MIA attacks increases as $r$ increases, showing less robustness to attacks for lager $r$. Right: The success rate of MIA increases as $r$ increases, showing less robustness to attacks for lager $r$.
We observed in Section 5 that the variance of the noise injected to the batch gradients w.r.t all existing adapters is inversely proportional to $r$. Therefore, we expect LoRA to show more robustness to MIA as $r$ decreases. This is observed in Figure 3, where we have done an ablation study on the Pubmed dataset about the effect of $r$. As observed in the left figure, the perplexity of the fine-tuned model increases as $r$ decreases, due to more noise injection into the adapters’ batch gradients during adaptation. Similarly, as observed in the middle and the right figure, robustness of the fine-tuned model w.r.t MIA increases as $r$ decreases. Again this can be attributed to the fact that larger noise will be added to batch gradients when we use smaller $r$. This interesting observation backed with our theoretical results shows that the adaptation rank $r$ in LoRA adjusts the trade-off between utility of the fine-tuned model and its robustness to MIA, just like the $\epsilon$ parameter for DPSGD.
Figure 4: Experimental results obtained on Pubmed dataset showing that, unlike DPSGD, LoRA adds noise to batch gradient w.r.t the adapter parameters non-uniformly. Left: the norm of the batch gradient w.r.t the adapters in different layers of the GPT-2 model. Batch gradient w.r.t the adapters located in the fifth and sixth layers have the largest norms. Right: The noise measure $N(W_{\ell}^{0},r)$ evaluated at different layers (indexed by $\ell$) of the model, showing that those layers with larger batch gradient w.r.t their adapter experience a large noise injection by LoRA. Also, as the adaptation rank $r$ increases the amount of noise injected to each layer decreases.
Now, we introduce a measure to quantify the amount of noise introduced by low-rank adaptation. Let $W_{\ell}^{0}$ denote the pretrained adapter parameter at the layer $\ell$ of the model. Also, let $W_{\ell}^{1}$ denote the adapter parameter that is obtained for the layer $\ell$ after one SGD update with step size $\eta$ w.r.t $W_{\ell}^{0}$. In that case, the forward pass of the adaptation layer $\ell$ will be updated to $y=W_{\ell}^{1}x$ (fig. 1). Similarly, let $A_{\ell}^{1}$ and $B_{\ell}^{1}$ denote the parameters obtained from low-rank adaptation of $W_{\ell}^{0}$ (with rank $r$) after one SGD update with the same step size $\eta$ w.r.t $A_{\ell}^{0}$ and $B_{\ell}^{0}$ (when matrix $A_{\ell}^{0}$ is frozen, we will have $A_{\ell}^{1}=A_{\ell}^{0}$). In that case, according to eq. 6, the forward pass of the adaptation layer $\ell$ will be updated to $y=(W_{\ell}^{0}+B_{\ell}^{1}A_{\ell}^{1})x$. According to our previous understandings that low-rank adaptation of $W_{\ell}^{0}$ is the same as fine-tuning it with noisy batch gradients, the amount of noise introduced by low-rank adaptation to the adapter in layer $\ell$ after one gradient step can be measured by:
$N(W_{\ell}^{0},r)=\big{\|}W_{\ell}^{1}-(W_{\ell}^{0}+B_{\ell}^{1}A_{\ell}^{1})\big{\|}.$ (21)
We have used the above noise measure in fig. 4 to confirm our theoretical findings that LoRA adds more noise to the adapter parameters that have a batch gradient with larger norm. We clearly observe that the adapter $W_{6}^{0}\in\mathbb{R}^{768\times 2304}$ in the sixth attention layer of GPT-2, which has the largest $\|\nabla_{W_{6}^{0}}\mathcal{L}\|$ (left figure), experiences the largest noise in one gradient step of LoRA (right figure). Also, as seen in the right figure, the amount of noise injected to different layers decreases uniformly as the rank $r$ increases. We already observed in fig. 3 that larger $r$ results in more vulnerability of the adapted model to privacy attacks, altogether confirming our theoretical findings in lemma 3.1 and section 5.
7 Conclusion
In this study, we uncovered a close similarity between the dynamics of low-rank adaptation and DPSGD. We showed that low-rank adaptation secretly injects random noise into the batch gradients w.r.t the adapter parameters non-uniformly. By quantifying the variance of this noise, we showed that the rank of adaptation controls the amount of injected noise. By bounding the distance between the injected noise distribution and pure Gaussian noise with the same variance, we demonstrated that the dynamics of low-rank adaptation closely imitates that of DPSGD. Backed with these theoretical results, we showed with our experimental results that low-rank adaptation indeed provides robustness against membership privacy attacks to the fine-tuning data. Compared to DPSGD, it has a less adverse effect on the utility of the fine-tuned model and a smaller run time. This suggests that low-rank adaptation can effectively balance utility of the fine-tuned model and its robustness to privacy attacks.
Acknowledgement
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Facebook award, and Google award. We also would like to thank Gautam Kamath (Cheriton School of Computer Science, University of Waterloo) and Abbas Mehrabian (Google DeepMind) for their valuable comments on some parts of this draft.
Appendix A Background
In this section, we provide some background about differential privacy and DPSGD algorithm, which are used in the next sections. In machine learning, differential privacy is often used to provide formal and rigorous data privacy guarantees w.r.t to training data. The following definition of differential privacy is one of the most commonly used definitions:
Definition A.1 (($\epsilon,\delta$)-DP, Dwork2006OurDO ).
A randomized mechanism $\mathcal{M}:\mathcal{D}\to\mathcal{R}$ with domain $\mathcal{D}$ and range $\mathcal{R}$ satisfies $(\epsilon,\delta)$-DP if for any two adjacent inputs $d$, $d^{\prime}\in\mathcal{D}$, which differ only by a single record, and for any measurable subset of outputs $\mathcal{S}\subseteq\mathcal{R}$ it holds that
$\texttt{Pr}[\mathcal{M}(d)\in\mathcal{S}]\leq e^{\epsilon}\texttt{Pr}[\mathcal{M}(d^{\prime})\in\mathcal{S}]+\delta.$
Gaussian mechanism, which randomizes the output of a non-private computation $f$ on a dataset $d$ as $\mathbf{G}_{\sigma}f(d)\triangleq f(d)+\mathcal{N}(\mathbf{0},\sigma^{2})$, provides ($\epsilon,\delta$)-DP. The variance of the noise, $\sigma^{2}$, is calibrated to the sensitivity of $f$, i.e. the maximum amount of change in its output (measured in $\ell_{2}$ norm) on two neighboring datasets $d$ and $d^{\prime}$. Gaussian mechanism has been used in DPSGD algorithm (abadi, ) for private ML by randomizing intermediate data-dependent computations, e.g. batch gradients. More specifically, at the $t$-th gradient update step on a current model $\bm{\theta}\in\mathbb{R}^{p}$, DPSGD computes the following noisy batch gradient:
$\tilde{g}(\bm{\theta})=\frac{1}{b}\bigg{[}\Big{(}\sum_{i\in\mathcal{B}^{t}}\bar{g}_{i}(\bm{\theta})\Big{)}+\mathcal{N}(\mathbf{0},\sigma_{\texttt{{DP}}}^{2}\mathbb{I}_{p})\bigg{]},$ (22)
where $\bar{g}_{i}(\bm{\theta})=\texttt{clip}(g_{i},c)$, $g_{i}$ is the sample gradient computed on the sample $i$ in the batch $\mathcal{B}^{t}$ and $c$ is a clipping threshold. For a given vector $\mathbf{v}$, $\texttt{clip}(\mathbf{v},c)=\min\{\|\mathbf{v}\|,c\}\cdot\frac{\mathbf{v}}{\|\mathbf{v}\|}$. Also, $\sigma_{\texttt{{DP}}}=c\cdot z$, where $z$ is the noise scale that should be used by DPSGD in order to achieve $(\epsilon,\delta)-$DP (for desired values of $\epsilon$ and $\delta$) with respect to train data at the end of training. It can be computed by using a privacy accountant, e.g. the moments accountant (abadi, ).
Appendix B Useful Theorems
In this section, we mention some theorems, which we will use in our proofs.
Theorem B.1 (Chi-Squared distribution: (AlexMood, ), Section 4.3, Theorem 7).
If the random variables $X_{i}$, $i=1,\ldots,k$, are normally and independently distributed with means $\mu_{i}$ and variances $\sigma_{i}^{2}$, then
$U=\sum_{i=1}^{k}\big{(}\frac{X_{i}-\mu_{i}}{\sigma_{i}}\big{)}^{2}$ (23)
has a chi-squared distribution with $k$ degrees of freedom: $U\sim\mathcal{X}_{k}^{2}$. Also, $\mathbb{E}[U]=k$ and $\texttt{Var}[U]=2k$.
The theorem above states that sum of the squares of $k$ standard normal random variables is a chi-squared distribution with $k$ degrees of freedom.
Lemma B.2 (Raw moment of Chi-Squared distribution).
Suppose $X\sim\mathcal{X}_{k}^{2}$. Then, the $m$-th raw moment of $X$ can be found as follows;
$\mathbb{E}[X^{m}]=\prod_{i=0}^{m-1}(k+2i)$ (24)
Proof.
From the definition of Chi-Squared distribution with $r$ degrees of reddom, $U$ has the following probability density function:
$f_{X}(x)=\frac{1}{2^{\frac{k}{2}}\Gamma(\frac{k}{2})}x^{\frac{k}{2}-1}e^{-\frac{x}{2}}$ (25)
Therefore, we have:
[…]
$=\prod_{i=0}^{m-1}(k+2i).$ (26)
Note that the fifth equality directly results from the property of gamma function that for $z>0$, $\Gamma(1+z)=z\Gamma(z)$. ∎
Theorem B.3 (Classical Central Limit Theorem: (patrickB, ), Theorem 27.1).
Suppose that $\{X_{i}\}_{i=1}^{n}$, is an independent sequence of random variables having the same distribution with mean $\mu$ and positive variance $\sigma^{2}$. Define $S_{n}=\sum_{i=1}^{n}X_{i}$ as their sum. Let $Z_{n}$ be defined by
$Z_{n}=\frac{S_{n}-n\mu}{\sqrt{n}\sigma}.$ (27)
Then, the distribution of $Z_{n}$ approaches standard normal distribution as $n$ approaches infinity.
The theorem above states that $S_{n}$ is approximately, or asymptotically, distributed as a normal distribution with mean $n\mu$ and variance $n\sigma^{2}$.
The next theorem is about the Lindeberg’s condition, which is a sufficient (and under certain conditions also a necessary condition) for the Central Limit Theorem (CLT) to hold for a sequence of independent random variables $\{X_{i}\}_{i=1}^{n}$. Unlike the classical CLT stated above, which requires the sequence of random variables to have a finite variance and be both independent and identically distributed (i.i.d), Lindeberg’s CLT only requires the sequence of random variables to have finite variance, be independent and also satisfy the Lindeberg’s condition. The following states the theorem.
Theorem B.4 (Lindeberg and Lyapounov Theorem: (patrickB, ), Theorem 27.2).
Suppose $X_{1},\ldots,X_{n}$ are $n$ independent random variables with $\mathbb{E}[X_{i}]=\mu_{i}$ and $\texttt{Var}[X_{i}]=\sigma_{i}^{2}>0$. Define $S_{n}=\sum_{i=1}^{n}X_{i}$ and let $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$. Also assume the following condition holds for all $\epsilon>0$:
[…]
where $P_{X_{i}}$ is the pdf of variable $X_{i}$. Assuming $Z_{n}=\frac{S_{n}-\sum_{i=1}^{n}\mu_{i}}{s_{n}}$, the distribution of $Z_{n}$ approaches standard normal distribution as $n$ approaches infinity.
The theorem above states that, given that Lindeberg’s condition is satisfied, $S_{n}$ is approximately, or asymptotically, distributed as a normal distribution with mean $\sum_{i=1}^{n}\mu_{i}$ and variance $s_{n}^{2}$, even if the sequence of variables are not identically distributed.
Appendix C Proofs
Using the theorems above, we are now able to prove lemma 3.1. See 3.1
Proof.
From the theorem’s assumption, we know that elements of $A$ are from $\mathcal{N}(0,\frac{1}{r})$. Therefore, we can rewrite the product $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ as the following product:
$q\cdot(\frac{A^{\top}A}{r}-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ (29)
where the elements of $A$ are now from standard normal distribution. Let $a_{i,j}$ denote the element in $i$-th row and $j$-th column of this new $A$. Therefore, for all $i$ and $j$, $a_{i,j}$ has distribution $\mathcal{N}(0,1)$. Let $B=\frac{A^{\top}A}{r}-\mathbb{I}_{m}$. Also, let $A_{i,:}$ and $A_{:,j}$ denote the $i$-th row and $j$-th column of the new $A$, respectively. We have:
$B_{i,i}=\frac{1}{r}[A^{\top}A]_{i,i}-1=\frac{1}{r}A_{:,i}^{\top}A_{:,i}-1=\frac{1}{r}\|A_{:,i}\|_{2}^{2}-1=(\frac{1}{r}\sum_{l=1}^{r}a_{l,i}^{2})-1$ (30)
From eq. 29, we know that $a_{l,i}$ is from standard normal distribution. Hence, $a_{l,i}^{2}$ is a chi-squared with 1 degree of freedom: $a_{l,i}^{2}\sim\mathcal{X}_{1}^{2}$. Therefore, $\sum_{l=1}^{r}a_{l,i}^{2}$, which is the sum of $r$ independent chi-squared variables with 1 degree of freedom, is a chi-squared with $r$ degrees of freedom: $\sum_{l=1}^{r}a_{l,i}^{2}\sim\mathcal{X}_{r}^{2}$ (see theorem B.1). Therefore, for $i\in\{1,\ldots,m\}$, we have:
$\mathbb{E}[B_{i,i}]=\mathbb{E}\big{[}\frac{\sum_{l=1}^{r}a_{l,i}^{2}}{r}\big{]}-1=\frac{r}{r}-1=0,$
$\texttt{Var}[B_{i,i}]=\texttt{Var}[\frac{\sum_{l=1}^{r}a_{l,i}^{2}}{r}]=\frac{\texttt{Var}(\mathcal{X}_{r}^{2})}{r^{2}}=\frac{2r}{r^{2}}=\frac{2}{r}.$ (31)
Similarly, we find the mean and variance of the non-diagonal elements $B_{i,j}(i\neq j)$ of $B$. We have:
$B_{i,j}=\frac{1}{r}[A^{\top}A]_{i,j}=\frac{1}{r}A_{:,i}^{\top}A_{:,j}=\frac{1}{r}\sum_{l=1}^{r}a_{l,i}a_{l,j},$ (32)
where $a_{l,i}$ and $a_{l,j}$ are independent and standard normal. Therefore, $a_{l,i}+a_{l,j}\sim\mathcal{N}(0,2)$. Similarly, $a_{l,i}-a_{l,j}\sim\mathcal{N}(0,2)$. So we can rewrite $a_{l,i}a_{l,j}$ as:
$a_{l,i}a_{l,j}=\frac{1}{4}(a_{l,i}+a_{l,j})^{2}-\frac{1}{4}(a_{l,i}-a_{l,j})^{2}=\frac{1}{2}z_{1}^{2}-\frac{1}{2}z_{2}^{2},$ (33)
where $z_{1}$ and $z_{2}$ are from standard normal. Therefore, $a_{l,i}a_{l,j}=\frac{\nu_{1}-\nu_{2}}{2}$, where $\nu_{1},\nu_{2}\sim\mathcal{X}_{1}^{2}$. Also, $a_{l,i}+a_{l,j}$ and $a_{l,i}-a_{l,j}$ are independent variables. Hence, $z_{1}$ and $z_{2}$ are independent, and likewise $\nu_{1}$ and $\nu_{2}$ are independent. We conclude that:
$a_{l,i}a_{l,j}=\frac{1}{2}(\nu_{1}-\nu_{2}),$ (34)
where $\nu_{1},\nu_{2}\sim\mathcal{X}_{1}^{2}$, and are independent.
Now, lets assume $\nu_{1},\nu_{2}\sim\mathcal{X}_{k}^{2}$ (a more general case), and let $M_{\nu_{1}}(t)=\mathbb{E}[e^{t\nu_{1}}]$ be the moment generating function (MGF) of $\nu_{1}$. In this case, we know that $M_{\nu_{1}}(t)=M_{\nu_{2}}(t)=(1-2t)^{-\frac{k}{2}}$ (MGF of $\mathcal{X}_{k}^{2}$). Hence, $M_{\nu_{1}-\nu_{2}}(t)=M_{\nu_{1}}(t)\cdot M_{\nu_{2}}(-t)=(1-4t^{2})^{-\frac{k}{2}}=\big{(}\frac{\frac{1}{4}}{\frac{1}{4}-t^{2}}\big{)}^{\frac{k}{2}}$, which is the MGF of a symmetric about origin variance-gamma distribution with parameters $\lambda=\frac{k}{2},\alpha=\frac{1}{2},\beta=0,\mu=0,\gamma=\frac{1}{2}$. Therefore, when $\nu_{1},\nu_{2}\sim\mathcal{X}_{k}^{2}$, then $\nu_{1}-\nu_{2}$ has this distribution, which has mean $\mu+2\beta\lambda/\gamma^{2}=0$ and variance $2\lambda(1+2\beta^{2}/\gamma^{2})/\gamma^{2}=4k$.
In eq. 34, we had $k=1$, as we had $\nu_{1},\nu_{2}\sim\mathcal{X}_{1}^{2}$. Hence, based on the discussion above, we have:
$\mathbb{E}[a_{l,i}a_{l,j}]$ $=0$ (35)
$\texttt{Var}[a_{l,i}a_{l,j}]$ $=\frac{1}{4}\texttt{Var}[\nu_{1}-\nu_{2}]=\frac{4k}{4}=1$ (36)
Consequently, based on eq. 32 and from the results above, we can compute the mean and variance of the non-diagonal elements of $B$ ($i\neq j$):
$\mathbb{E}[B_{i,j}]=\mathbb{E}\big{[}\frac{\sum_{l=1}^{r}a_{l,i}a_{l,j}}{r}\big{]}=\frac{\sum_{l=1}^{r}\mathbb{E}[a_{l,i}a_{l,j}]}{r}=0,$
$\texttt{Var}[B_{i,j}]=\texttt{Var}[\frac{\sum_{l=1}^{r}a_{l,i}a_{l,j}}{r}]=\frac{\sum_{l=1}^{r}\texttt{Var}[a_{l,i}a_{l,j}]}{r^{2}}=\frac{r}{r^{2}}=\frac{1}{r}.$ (37)
So far, we have computed the mean and variance of each entry in $B=\frac{A^{\top}A}{r}-\mathbb{I}_{m}\in\mathbb{R}^{m\times m}$ in appendix C and appendix C. Now, for a given $q\in\mathbb{R}^{1\times m}$, we have:
$q\cdot B=\sum_{l=1}^{m}q_{l}B_{l,:},$ (38)
where $B_{l,:}$ is row $l$ of $B$. Let $u_{i}$ denote the $i$-th element of $q\cdot B$. Hence, for each element $u_{i}$ ($i\in\{1,\ldots,m\}$), we have:
$\mathbb{E}[u_{i}]$ $=\mathbb{E}\bigg{[}\sum_{l=1}^{m}q_{l}B_{l,i}\bigg{]}=\sum_{l=1}^{m}q_{l}\mathbb{E}[B_{l,i}]=0,$
$\texttt{Var}[u_{i}]$ $=\texttt{Var}\bigg{[}\sum_{l=1}^{m}q_{l}B_{l,i}\bigg{]}=\sum_{l=1}^{m}q_{l}^{2}\texttt{Var}[B_{l,i}]=q_{i}^{2}\texttt{Var}[B_{i,i}]+\sum_{l\neq i}q_{l}^{2}\texttt{Var}[B_{l,i}]$
$=q_{i}^{2}\frac{2}{r}+\sum_{l\neq i}q_{l}^{2}\frac{1}{r}=\frac{q_{i}^{2}}{r}+\sum_{l=1}^{m}q_{l}^{2}\frac{1}{r}=\frac{q_{i}^{2}+\sum_{l=1}^{m}q_{l}^{2}}{r}\approx\frac{\sum_{l=1}^{m}q_{l}^{2}}{r}=\frac{\|q\|_{2}^{2}}{r},$ (39)
where the approximation is indeed valid because $m$, which is the dimension of the input of the current layer (see fig. 1), is a large integer. Finally, according to eq. 38, each element $u_{i}$ of $qB$ is the sum of $m$ random variables, for which the Lindeberg’s condition is also satisfied: as $m\to\infty$, $s_{m}^{2}=\frac{\|q\|_{2}^{2}}{r}\to\infty$ ($m$ is the dimension of $q$, and $s_{m}$ is the sum of variances of the $m$ random variables, which we found in appendix C). Hence, $[|u_{i}-0|>\epsilon s_{m}]\downarrow\emptyset$ as $m\to\infty$. Therefore, from theorem B.4, we also conclude that as $m\to\infty$, each element of $qB$ approaches a Gaussian with the mean and variance found in appendix C. Therefore, we conclude that having an $A$, where the elements of $A$ are i.i.d and from $\mathcal{N}(0,\frac{1}{r})$, then as $m\to\infty$, $q\cdot(A^{\top}A-\mathbb{I}_{m})\in\mathbb{R}^{1\times m}$ approaches a Gaussian $\mathcal{N}(0,\frac{\|q\|^{2}}{r})$, which completes the proof.
See 4.2
Proof.
From eq. 38, we had:
$u_{i}=\sum_{l\neq i,l=1}^{m}q_{l}B_{l,i}+q_{i}B_{i,i},$ (40)
where $B_{l,i}=\frac{1}{r}A^{\top}_{:,l}A_{:,i}=\frac{1}{2r}\sum_{t=1}^{r}V_{t}$, where $V_{t}\sim\texttt{Variance-Gamma}(\nu,\alpha,\beta,\mu)$ with $\nu=\beta=\mu=0$ and $\alpha=\frac{1}{2}$. Also $B_{i,i}=\frac{1}{r}A^{\top}_{:,i}A_{:,i}-1=\frac{\texttt{X}}{r}-1$, where $X\sim\mathcal{X}_{r}^{2}$. Therefore, we can rewrite the equation above for $u_{i}$ as:
$u_{i}=\sum_{l\neq i,l=1}^{m}\frac{q_{l}}{2r}\sum_{t=1}^{r}V_{t}+q_{i}(\frac{\texttt{X}}{r}-1)=\sum_{l\neq i,l=1}^{m}\sum_{t=1}^{r}\frac{q_{l}}{2r}V_{t}+\frac{q_{i}}{r}(\texttt{X}-r),$ (41)
where $V_{t}\sim\texttt{Variance-Gamma}(\nu,\alpha,\beta,\mu)$ with $\nu=\beta=\mu=0$ and $\alpha=\frac{1}{2}$ and $X\sim\mathcal{X}_{r}^{2}$. Hence, $V_{t}$ has mean 00 and variance $4$ and $(\texttt{X}-r)$ has mean 00 and variance $2r$. Also note that $X$ can be written as the summation of $r$ independent variables with distribution $\mathcal{X}_{1}^{2}$. Therefore, $u_{i}$ is the weighted sum of $mr$ independent random variables with mean 00. Also, from appendix C in the proof of lemma 3.1, we know that $u_{i}$ has mean 00 and variance $\frac{\|q\|_{2}^{2}}{r}$. Now, in order to bound the TV distance between the distribution of $u_{i}$ and $\mathcal{N}(0,\frac{\|q\|_{2}^{2}}{r})$, we have to use theorem 4.1 and eq. 11. More specifically, we have to find the third Lyapounov ratio $L_{3}=\frac{\sum_{i}\mathbb{E}[|X_{i}|^{3}]}{s_{n}^{3}}=\frac{\sum_{i}\mathbb{E}[|X_{i}|^{3}]}{\big{(}\sum_{i}\texttt{Var}[X_{i}]\big{)}^{3}}=\frac{\sum_{i}\mathbb{E}[|X_{i}|^{3}]}{\big{(}\sum_{i}\mathbb{E}[X_{i}^{2}]\big{)}^{3}}$, where $X_{i}$ is each of the $1+(m-1)r$ summands in eq. 41. First we note that, based on appendix C, $s_{n}^{3}=(\frac{\|q\|_{2}^{2}}{r})^{\frac{3}{2}}=\frac{\|q\|_{2}^{3}}{r\sqrt{r}}$. Now, we find the numerator $\sum_{i}\mathbb{E}[|X_{i}|^{3}]$. From (gaunt2024, ), we know that for $V_{t}\sim\texttt{Variance-Gamma}(\nu,\alpha,0,0),\mathbb{E}[|V_{t}|^{r}]=\frac{2^{r}}{\sqrt{\pi}\alpha^{r}}\frac{\Gamma(\nu+(r+1)/2)\Gamma((r+1)/2)}{\Gamma(\nu+1/2)}$. Therefore, for $V_{t}\sim\texttt{Variance-Gamma}(0,\frac{1}{2},0,0),\mathbb{E}[|V_{t}|^{3}]=\frac{2^{6}}{\pi}$. On the other hand, we know that the skewness of $X\sim\mathcal{X}_{r}^{2}$ is equal to $\frac{\mathbb{E}[(X-\mathbb{E}[X])^{3}]}{\texttt{Var}[X]^{\frac{3}{2}}}=\frac{\mathbb{E}[(X-r)^{3}]}{(2r)^{\frac{3}{2}}}=\sqrt{\frac{8}{r}}$. Hence, $\mathbb{E}[(X-r)^{3}]=(2r)^{\frac{3}{2}}\sqrt{\frac{8}{r}}=8r$. Hence for $X\sim\mathcal{X}_{r}^{2}$, $\mathbb{E}[|X-r|^{3}]\geq\mathbb{E}[(X-r)^{3}]=8r$. Now, we can find the numerator $\sum_{i}\mathbb{E}[|X_{i}|^{3}]$ as:
$\sum_{i}\mathbb{E}[|X_{i}|^{3}]$ $=\sum_{l\neq i,l=1}^{m}\sum_{t=1}^{r}\frac{|q_{l}|^{3}}{8r^{3}}\mathbb{E}[|V_{t}|^{3}]+\frac{|q_{i}|^{3}}{r^{3}}\mathbb{E}[|\texttt{X}-r|^{3}]$
$=\sum_{l\neq i,l=1}^{m}\frac{|q_{l}|^{3}}{8r^{2}}\cdot\frac{2^{6}}{\pi}+\frac{|q_{i}|^{3}}{r^{3}}\mathbb{E}[|\texttt{X}-r|^{3}]$
$\approx\sum_{l\neq i,l=1}^{m}\frac{8|q_{l}|^{3}}{\pi r^{2}}+\frac{8|q_{i}|^{3}}{r^{2}}\approx\sum_{l=1}^{m}\frac{8|q_{l}|^{3}}{\pi r^{2}}=\frac{8}{\pi r^{2}}\|q\|_{3}^{3}.$ (42)
Therefore, for the sum $u_{i}$ in eq. 41, we have the third Lyapounov ratio:
$L_{3}=\frac{8}{\pi r^{2}}\|q\|_{3}^{3}\times\frac{r\sqrt{r}}{\|q\|_{2}^{3}}=\frac{8}{\pi\sqrt{r}}\bigg{(}\frac{\|q\|_{3}}{\|q\|_{2}}\bigg{)}^{3}.$ (43)
Therefore, based on theorem 4.1, we have:
$\|Q_{m}(x)-\Phi(x)\|_{TV}\leq\frac{8C_{D}}{\pi\sqrt{r}}\bigg{(}\frac{\|q\|_{3}}{\|q\|_{2}}\bigg{)}^{3},$ (44)
where $C_{D}\leq\frac{\pi\sqrt{r}}{8}$ is a constant, which depends only on $D$, where $D$ is an upperbound for the KL divergence between each of the random variable summands in eq. 41 and a Gaussian with the same mean and variance. Now, assuming $0<c\leq|q_{i}|\leq C$ for the elements $q_{i}$ in $q$, we have $\bigg{(}\frac{\|q\|_{3}}{\|q\|_{2}}\bigg{)}^{3}\leq\bigg{(}\frac{|C|}{|c|}\bigg{)}^{3}\frac{1}{\sqrt{m}}$. Therefore:
$\|Q_{m}(x)-\Phi(x)\|_{TV}\leq\frac{8C_{D}}{\pi}\bigg{(}\frac{|C|}{|c|}\bigg{)}^{3}\frac{1}{\sqrt{mr}}.$ (45)
Therefore,
$\|Q_{m}(x)-\Phi(x)\|_{TV}\in\mathcal{O}\bigg{(}\frac{1}{\sqrt{mr}}\bigg{)}.$ (46)